### v1

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract
Representing vector fields by scalar or vector potentials can be a challenging task in domains with cavities or tunnels, due to the presence of harmonic fields which are both irrotational and solenoidal but may have no scalar or vector potentials. For harmonic fields normal to the boundary, which can exist in domains with cavities, it is possible to define scalar potentials with Dirichlet boundary conditions fitted to the domain’s cavities. For harmonic fields tangent to the boundary, which can exist in domains with tunnels, a similar construction was lacking. In this article we present a construction of vector potentials that yield a basis for the tangent harmonic fields. Our vector potentials are obtained by solving curl-curl problems with inhomogeneous tangent boundary conditions that are fitted to closed curves looping around the tunnels. Applied to structure-preserving finite elements, our method also provides an exact geometric parametrization of the discrete harmonic fields.
1 Introduction
Using scalar or vector potentials to represent vector fields is a powerful technique in physics and engineering. In practical computations they ensure that a vector field is either irrotational (curl-free) or solenoidal (divergence-free), which has useful applications in many branches such as astrophysics [35, 34], magnetostatics [7, 32] or fluid dynamics [28], and they are also involved in multigrid solvers [3, 23]. Beyond numerical applications, potentials are fundamental to several variational formulations, for instance in kinetic or fluid electromagnetic models [25, 27] and they play a crucial role in quantum [17] or relativistic physics [33]: a recent illustration is the elegant reformulation of the relativistic Vlasov-Maxwell equations as a wave-transport system involving Liénard-Wiechert potentials [11, 12]. Finally, their geometric properties expressed in Stokes’ theorem relate to key topological aspects of physical laws, in particular in electromagnetism [10, 19].
A difficulty arises when considering domains with cavities or tunnels, where harmonic fields with vanishing traces may exist [31, 2]. Harmonic fields are unique because they are both irrotational and solenoidal, yet they may not admit scalar or vector potentials depending on their boundary conditions. Specifically, domains with cavities have harmonic fields normal to the boundary that admit scalar potentials but no vector potentials, while domains with tunnels have harmonic fields tangent to the boundary that admit vector potentials but no scalar potentials.
These cavities and tunnels correspond to the 2-chain and 1-chain homology groups of the domain, whose ranks determine the dimensions of the spaces of normal and tangent harmonic vector fields, respectively [24, 2]: in particular these dimensions are topological invariants, which means that they remain unchanged under continuous deformations (with continuous inverse) of the domain. For normal harmonic fields, a construction has been proposed for scalar potentials, characterized by Laplace equations with boundary conditions fitted to the domain’s cavities [31]. For tangent harmonic fields however, a similar construction is more complex: in domains diffeomorphic to a solid torus, a common approach is to use the toroidal angle as a (piecewise smooth) scalar potential [26, 20], and this approach has been extended to a large class of domains, namely domains that can be made simply connected by removing a finite number of disjoint cuts surfaces [18, 31, 1]. A construction of proper vector potentials, however, seems to be lacking for tangent harmonic fields.
In this article we propose such a construction. Our vector potentials are characterized by homogeneous curl-curl problems, with specific tangent boundary conditions fitted to closed curves that loop around the tunnels of the domain. These “tunnel curves” form a generating family for the 1-chain homology group, and the associated vector potentials yield a basis for the space of tangent harmonic vector fields. In our analysis we assume the existence of surfaces reciprocal to the tunnel curves: these surfaces are reminiscent of the cut surfaces involved in the construction of piecewise smooth scalar potentials, but here their purpose is only to establish the linear independence of the resulting harmonic fields. In particular they do not need not be disjoint, which allows us to apply our construction to domains such as hollow tori.
An interesting feature of our approach is that it can also be applied in a structure-preserving finite element framework. Under some assumptions which are fulfilled by Whitney [36, 9] and higher order Nédélec elements [29, 30], and more generally by the discrete de Rham sequences constructed by Hiptmair [21] and Arnold, Falk and Winther [4, 5] in the Finite Element Exterior Calculus (FEEC) framework, a natural discretization of our construction leads to an exact geometric parametrization of the discrete harmonic fields.
The article is organized as follows. In Section 2 we begin by reminding the definition of the homology groups, together with that of the harmonic spaces in the $L^{2}$ de Rham complex, and we discuss some important properties regarding their dimensions and potentials. In Section 3 we then describe our construction, which is based on a decomposition of the vector potentials in two parts: a smooth part that lifts a specific boundary condition fitted to the tunnel curves, and a correction part that guarantees that the resulting potential solves an homogeneous curl-curl problem. Section 4 is finally devoted to studying a discrete version of our construction. After listing some properties that several structure-preserving finite element frameworks possess, and discussing the case of the discrete normal harmonic fields, we show that a natural application of our approach yields a basis of the discrete tangent harmonic fields, in terms of discrete vector potentials associated to the tunnels of the domain.
In our work, an important source of inspiration has been the geometric approach promoted since the 1980s by Alain Bossavit for the development of structure-preserving methods in computational electromagnetism [9, 10, 32]. For instance, a construction of discrete vector potentials for magnetic fields satisfying specific flux conditions is described in [10], within the framework of Whitney elements. We learned of Alain Bossavit’s passing as we were completing this work, which we dedicate to his memory.
2 Homology groups and harmonic function spaces
Throughout the article, we consider that $\Omega$ is an open bounded domain in $\mathbb{R}^{3}$ with a Lipschitz boundary, in the usual sense that $\partial\Omega$ is locally the graph of Lipschitz mappings, with $\Omega$ on one side only. For simplicity, we assume that $\Omega$ is connected.
2.1 Harmonic fields in the Hilbert de Rham complex
Let us start by reminding how the Hilbert spaces of harmonic functions are defined in $\Omega$: one first considers the Hilbert de Rham complex [10, 5]
$0\xrightarrow{~~}H^{1}(\Omega)\xrightarrow{~\operatorname{{\bf grad}}~}H(\operatorname{{\bf curl}};\Omega)\xrightarrow{~\operatorname{{\bf curl}}~}H(\operatorname{div};\Omega)\xrightarrow{~\operatorname{div}~}L^{2}(\Omega)\xrightarrow{~~}0$ (1)
for which the dual complex, made of the $L^{2}$ adjoint operators and their respective domains, is
$0\xleftarrow{~~}L^{2}(\Omega)\xleftarrow{~-\operatorname{div}~}H_{0}(\operatorname{div};\Omega)\xleftarrow{~\operatorname{{\bf curl}}~}H_{0}(\operatorname{{\bf curl}};\Omega)\xleftarrow{~-\operatorname{{\bf grad}}~}H^{1}_{0}(\Omega)\xleftarrow{~~}0.$ (2)
Following [4], we rewrite the primal complex using a generic notation
$0\xrightarrow{~~}V^{0}\xrightarrow{~d^{0}~}V^{1}\xrightarrow{~d^{1}~}V^{2}\xrightarrow{~d^{2}~}V^{3}\xrightarrow{~~}0$ (3)
and we denote the kernel and range spaces by
$\mathfrak{Z}^{k}:=\{v\in V^{k}:d^{k}v=0\},\qquad\mathfrak{B}^{k}:=d^{k-1}V^{k-1}.$
The dual complex is denoted as
$0\xleftarrow{~~}V^{*}_{0}\xleftarrow{~d^{*}_{1}~}V^{*}_{1}\xleftarrow{~d^{*}_{2}~}V^{*}_{2}\xleftarrow{~d^{*}_{3}~}V^{*}_{3}\xleftarrow{~~}0$ (4)
with similar notation for the adjoint kernel and range spaces, i.e., $\mathfrak{Z}^{*}_{k}:=\{v\in V^{*}_{k}:d^{*}_{k}v=0\}$ and $\mathfrak{B}^{*}_{k}:=d^{*}_{k+1}V^{*}_{k+1}$. The harmonic spaces are then formally defined as
$\mathfrak{H}^{k}:=\mathfrak{Z}^{k}\cap(\mathfrak{B}^{k})^{\perp}=\mathfrak{Z}^{k}\cap\mathfrak{Z}^{*}_{k}$ (5)
which yields for the Hilbert de Rham complex above:
$\left\{\begin{aligned} &\mathfrak{H}^{0}=\mathbb{R},\\ &\mathfrak{H}^{1}=H(\operatorname{{\bf curl}}0;\Omega)\cap H_{0}(\operatorname{div}0;\Omega),\\ &\mathfrak{H}^{2}=H(\operatorname{div}0;\Omega)\cap H_{0}(\operatorname{{\bf curl}}0;\Omega),\\ &\mathfrak{H}^{3}=\{0\}\end{aligned}\right$. (6)
where we have used standard concise notation for the kernel spaces of the curl and divergence operators with or without boundary conditions, see e.g. [6]. If the domain is contractible (i.e., simply connected with a connected boundary), the above spaces of harmonic vector fields must vanish as a consequence of the Poincaré lemma [10]. In a general domain however this is not the case.
2.2 Singular chains and homology groups
The works of de Rham [16] have established profound connections between harmonic spaces and homology groups. Here we will follow [24] and consider singular homology groups. We remind that a singular $k$-simplex is a continuous map $\sigma:\Delta_{k}\to\Omega$ defined on the standard $k$-simplex $\Delta_{k}$, and that a (singular) $k$-chain is a formal linear combination of such singular simplices, with coefficients in $\mathbb{Z}$. Denoting by $\mathcal{C}_{k}(\Omega)$ the corresponding group of $k$-chains, and by $\partial:\mathcal{C}_{k}(\Omega)\to\mathcal{C}_{k-1}(\Omega)$ the (singular) boundary operator, we obtain a chain complex by considering the sequence
$0\xleftarrow{~~}\mathcal{C}_{0}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{1}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{2}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{3}(\Omega)\xleftarrow{~~}0.$ (7)
Here, the order from right to left corresponds to the orientation of the dual complex (2). Denoting by $\mathcal{Z}_{k}$ and $\mathcal{B}_{k}$ the $k$-cycles and $k$-boundaries, which are the kernels and ranges (in $\mathcal{C}_{k}(\Omega)$) of the boundary operator, and observing that the inclusion $\mathcal{B}_{k}\subseteq\mathcal{Z}_{k}$ follows from the sequence property $\partial\circ\partial=0$, one next defines the $k$-th homology group as the quotient
$\mathcal{H}_{k}:=\mathcal{Z}_{k}\operatorname{/}\mathcal{B}_{k}.$ (8)
Thus, an element of $\mathcal{H}_{k}$ is a class of closed $k$-chains (closed in a manifold sense, not a topological one) modulo boundary chains. Intuitively, this corresponds to the set of points, closed curves or closed surfaces (there are no closed volumes in $\Omega$) that can be obtained from each other by continuous deformations within the domain $\Omega$. An illustration is provided in Figure 1.
$\Omega$
[…]
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\Gamma}_{1,1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\p$gfsys@color@gray@fill{0}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\Gamma}_{1,2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\p$gfsys@color@gray@fill{0}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\Gamma}_{2,1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\p$gfsys@color@gray@fill{0}
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\tilde{\Gamma}_{2,2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\p$gfsys@color@gray@fill{0}
Figure 1: Example of a non-contractible domain $\Omega$ with two cavities ($\beta_{2}=2$) and two tunnels ($\beta_{1}=2$): one that goes through the domain, and another one that is also a cavity and circles around the first tunnel. The domain boundary has three connected components which are closed surfaces (in orange, only partially drawn for the sake of visibility): $S_{0}$ is the “outer” boundary which bounds the unbounded component of $\mathbb{R}^{3}\setminus\Omega$, $S_{1}$ bounds the contractible cavity, and $S_{2}$ bounds the inner tunnel. The last two represent independent classes of the 2-chain homology group $\mathcal{H}_{2}$, while the first one is homologous to $(S_{1}+S_{2})$. Finally, the closed curves $\Gamma_{i}\subset\partial\Omega$, $i=1,2$, (in light blue) represent two independent classes of the 1-chain homology group $\mathcal{H}_{1}$. They admit reciprocal surfaces $\Sigma_{i}$ as described in Assumption 3.1, visualized by their boundary curves $\tilde{\Gamma}_{i,a}\subset\partial\Omega$, drawn in red.
The rank of the $k$-th homology group is called the $k$-th Betti number of $\Omega$, and denoted
$\beta_{k}:=\dim(\mathcal{H}_{k}).$ (9)
The extremal numbers are simple: we always have $\beta_{3}=0$ since there are no closed volumes (three-dimensional cycles) in $\Omega$, and $\beta_{0}$ is the number of connected components, hence $\beta_{0}=1$ for a connected domain. The intermediate Betti numbers are more interesting and characterize the lack of contractibility of $\Omega$. In particular, $\beta_{2}$ is the number of cavities in the domain, i.e., the number of connected components of $\mathbb{R}^{3}\setminus\Omega$ which are bounded, and $\beta_{1}$ is the number of tunnels. Notice that tunnels may go through $\Omega$ like handles, or be entirely surrounded by it, which makes them cavities as well: an example is shown in Figure 1.
In the sequel we will denote by $S_{i}$, $i=0,\dots,\beta_{2}$, the connected components of the boundary $\partial\Omega$ (oriented by the normal vector pointing outside of $\Omega$). They are all closed surfaces: we let $S_{0}$ be the boundary of the single unbounded connected component of $\mathbb{R}^{3}\setminus\Omega$, so that for every $i=1,\dots,\beta_{2}$, $S_{i}$ bounds a single cavity surrounded by the domain $\Omega$: these surfaces form a basis of the 2-chain homology group $\mathcal{H}_{2}$. Observe that $S_{0}$ is homologous to $(\sum_{i=1}^{\beta_{2}}S_{i})$ since the difference bounds the domain itself.
Similarly, we will denote by $\Gamma_{i}$, $i=1,\dots,\beta_{1}$, a collection of oriented closed curves that generate the 1-chain homology group $\mathcal{H}_{1}$. These curves will be chosen on the boundary $\partial\Omega$, such that each one circles once around a different tunnel: a precise description will be discussed in Section 3.1 below.
2.3 The de Rham theorem for the $L^{2}$ de Rham complex
An important result [4, 2], which essentially follows from de Rham’s theorem [15, 13, 24], is that the dimensions of the harmonic spaces (5) coincide with the Betti numbers of the domain.
Theorem 2.1.
For $k=0,\dots,3$, the following relation holds:
$\dim(\mathfrak{H}^{k})=\beta_{k}.$ (10)
Since the classical version of de Rham’s theorem is formulated for the smooth complex, we give here a complete argument which uses a result of Costabel and McIntosh [14], suggested to us by Douglas Arnold.
Proof.
As an open subset of $\mathbb{R}^{3}$, $\Omega$ is a smooth manifold and the de Rham theorem applies, for instance as formulated in [24, Th. 18.14]. As a consequence, the $k$-chain cohomology groups (isomorphic to the dual spaces $(\mathcal{H}_{k})^{\prime}$ and hence of the same dimensions $\beta_{k}$) are isomorphic to the cohomology groups of the smooth de Rham complex
$0\xrightarrow{~~}C^{\infty}(\Omega)\xrightarrow{~\operatorname{{\bf grad}}~}C^{\infty}(\Omega)^{3}\xrightarrow{~\operatorname{{\bf curl}}~}C^{\infty}(\Omega)^{3}\xrightarrow{~\operatorname{div}~}C^{\infty}(\Omega)\xrightarrow{~~}0.$ (11)
Rewriting the latter as $\dots\xrightarrow{}V^{k}_{\infty}\xrightarrow{~d^{k}~}V^{k+1}_{\infty}\xrightarrow{}\dots$ and denoting by $\mathfrak{Z}^{k}_{\infty}:=\{v\in V^{k}_{\infty}:d^{k}v=0\}$ and $\mathfrak{B}^{k}_{\infty}:=d^{k-1}V^{k-1}_{\infty}$ the associated kernel and range spaces, one has then
$\dim\big{(}\mathfrak{Z}^{k}_{\infty}\operatorname{/}\mathfrak{B}^{k}_{\infty}\big{)}=\beta_{k}.$ (12)
Moreover, a Poincaré duality theorem holds ([24, Pbm. 18.7]), which states that the cohomology groups of the compactly supported de Rham sequence (indexed with reverse order) have the same dimensions. Namely, writing
$0\xleftarrow{~~}C^{\infty}_{c}(\Omega)\xleftarrow{~\operatorname{div}~}C^{\infty}_{c}(\Omega)^{3}\xleftarrow{~\operatorname{{\bf curl}}~}C^{\infty}_{c}(\Omega)^{3}\xleftarrow{~\operatorname{{\bf grad}}~}C^{\infty}_{c}(\Omega)\xleftarrow{~~}0$ (13)
as $\dots\xleftarrow{}V^{\infty}_{c,k-1}\xleftarrow{~d^{*}_{k}~}V^{\infty}_{c,k}\xleftarrow{}\dots$ and denoting the associated kernel and range spaces by $\mathfrak{Z}^{\infty}_{c,k}:=\{v\in V^{\infty}_{c,k}:d^{*}_{k}v=0\}$ and $\mathfrak{B}^{\infty}_{c,k}:=d^{*}_{k+1}V^{\infty}_{c,k+1}$, one has
$\dim\big{(}\mathfrak{Z}^{\infty}_{c,k}\operatorname{/}\mathfrak{B}^{\infty}_{c,k}\big{)}=\dim\big{(}\mathfrak{Z}^{k}_{\infty}\operatorname{/}\mathfrak{B}^{k}_{\infty}\big{)}.$ (14)
To link the above relations with the Hilbert harmonic spaces (5), we next invoke two results from [14]. The first one is Theorem 1.1, which states that the cohomology spaces of the sequences
$H^{-1}_{\overline{\Omega}}(\mathbb{R}^{3})\xleftarrow{~d^{*}_{k}~}H^{0}_{\overline{\Omega}}(\mathbb{R}^{3})\xleftarrow{~d^{*}_{k+1}~}H^{1}_{\overline{\Omega}}(\mathbb{R}^{3})\qquad\text{ and }\qquad V^{\infty}_{c,k-1}\xleftarrow{~d^{*}_{k}~}V^{\infty}_{c,k}\xleftarrow{~d^{*}_{k+1}~}V^{\infty}_{c,k+1}$ (15)
can be represented with the same functions. Here, $H^{s}_{\overline{\Omega}}(\mathbb{R}^{3})$ is the space of distributions in $H^{s}(\mathbb{R}^{3})$ with support in $\overline{\Omega}$, and the differential operators are meant in the sense of distributions over $\mathbb{R}^{3}$. We thus have
$\dim\big{(}\mathfrak{Z}^{*}_{\overline{\Omega},k}\operatorname{/}\mathfrak{B}^{*}_{\overline{\Omega},k}\big{)}=\dim\big{(}\mathfrak{Z}^{\infty}_{c,k}\operatorname{/}\mathfrak{B}^{\infty}_{c,k}\big{)}$ (16)
with
$\mathfrak{Z}^{*}_{\overline{\Omega},k}:=\{v\in L^{2}_{\overline{\Omega}}(\mathbb{R}^{3}):d^{*}_{k}v=0\}\qquad\text{ and }\qquad\mathfrak{B}^{*}_{\overline{\Omega},k}:=d^{*}_{k+1}H^{1}_{\overline{\Omega}}(\mathbb{R}^{3}).$
The second useful result from [14] is Corollary 4.7, which allows to write
$d^{*}_{k+1}H^{1}_{\overline{\Omega}}(\mathbb{R}^{n})=d^{*}_{k+1}H_{\overline{\Omega}}(d^{*}_{k+1};\mathbb{R}^{n})$
where the last space is defined as $H_{\overline{\Omega}}(d^{*}_{k+1};\mathbb{R}^{n}):=\{v\in L^{2}_{\overline{\Omega}}(\mathbb{R}^{3}):d^{*}_{k+1}v\in L^{2}_{\overline{\Omega}}(\mathbb{R}^{3})\}$. Since this space consists of the $L^{2}(\mathbb{R}^{3})$ functions which have a differential in $L^{2}$ and vanish outside of $\Omega$, it coincides with the space $V^{*}_{k+1}$ from the complex with homogeneous boundary conditions (2). As a result, we find
$\mathfrak{B}^{*}_{\overline{\Omega},k}=d^{*}_{k+1}H^{1}_{\overline{\Omega}}(\mathbb{R}^{3})=d^{*}_{k+1}V^{*}_{k+1}=\mathfrak{B}^{*}_{k},$
and we also observe that
$\mathfrak{Z}^{*}_{\overline{\Omega},k}=\{v\in H_{\overline{\Omega}}(d^{*}_{k};\mathbb{R}^{n}):d^{*}_{k}v=0\}=\{v\in V^{*}_{k}:d^{*}_{k}v=0\}=\mathfrak{Z}^{*}_{k}.$
It follows that (16) rewrites as
$\dim\big{(}\mathfrak{Z}^{*}_{k}\operatorname{/}\mathfrak{B}^{*}_{k}\big{)}=\dim\big{(}\mathfrak{Z}^{\infty}_{c,k}\operatorname{/}\mathfrak{B}^{\infty}_{c,k}\big{)}.$ (17)
The proof is finally completed by observing that the orthogonal projection onto $(\mathfrak{B}^{*}_{k})^{\perp}$ induces an isomorphism between the quotient space $\mathfrak{Z}^{*}_{k}\operatorname{/}\mathfrak{B}^{*}_{k}$ and the harmonic space $\mathfrak{H}^{k}=\mathfrak{Z}^{*}_{k}\cap(\mathfrak{B}^{*}_{k})^{\perp}$. ∎
2.4 Scalar and vector potentials
An important property of the complexes (1) and (2) is that the range spaces $\mathfrak{B}^{k}$ and $\mathfrak{B}^{*}_{k}$ are closed in $L^{2}(\Omega)$ [4]. As a result, one can write a general Hodge-Helmholtz decomposition of the form
$L^{2}(\Omega)=\mathfrak{B}^{k}\overset{\perp}{\oplus}\mathfrak{H}^{k}\overset{\perp}{\oplus}\mathfrak{B}^{*}_{k}$ (18)
which yields two $L^{2}$-orthogonal decompositions for vector fields: one that corresponds to $k=1$,
$L^{2}(\Omega)=\operatorname{{\bf grad}}H^{1}(\Omega)\overset{\perp}{\oplus}\mathfrak{H}^{1}\overset{\perp}{\oplus}\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$ (19)
and another one corresponding to $k=2$,
$L^{2}(\Omega)=\operatorname{{\bf curl}}H(\operatorname{{\bf curl}};\Omega)\overset{\perp}{\oplus}\mathfrak{H}^{2}\overset{\perp}{\oplus}\operatorname{{\bf grad}}H^{1}_{0}(\Omega).$ (20)
With regard to potentials, the question is then to know whether harmonic vector fields belong to some range space. It is clear from (19) that the tangential harmonic fields in $\mathfrak{H}^{1}$ have no scalar potential, from (20) we see that the normal harmonic fields in $\mathfrak{H}^{2}$ have no vector potential. But the reverse may be true, and a simple argument shows that it is indeed.
Proposition 2.1.
The following inclusions hold:
$H_{0}(\operatorname{{\bf curl}}0;\Omega)\subseteq\operatorname{{\bf grad}}H^{1}(\Omega),$ (21)
and
$H_{0}(\operatorname{div}0;\Omega)\subseteq\operatorname{{\bf curl}}H(\operatorname{{\bf curl}};\Omega).$ (22)
In particular, every field in $\mathfrak{H}^{1}$ admits a vector potential in $H(\operatorname{{\bf curl}};\Omega)$, and every field in $\mathfrak{H}^{2}$ admits a vector potential in $H^{1}(\Omega)$.
Proof.
If ${\bm{v}}\in H_{0}(\operatorname{{\bf curl}}0;\Omega)$, resp. ${\bm{v}}\in H_{0}(\operatorname{div}0;\Omega)$, then letting $\bar{\bm{v}}$ be its extension by 0 on an open ball $B_{R}$ that contains $\overline{\Omega}$ yields $\bar{\bm{v}}\in H_{0}(\operatorname{{\bf curl}}0;B_{R})$, resp. $\bar{\bm{v}}\in H_{0}(\operatorname{div}0;B_{R})$. Since $B_{R}$ is contractible this allows to find $\bar{\phi}\in H^{1}_{0}(B_{R})$ such that $\bar{\bm{v}}=\operatorname{{\bf grad}}\bar{\phi}$, resp. $\bar{\bm{A}}\in H_{0}(\operatorname{{\bf curl}};B_{R})$ such that $\bar{\bm{v}}=\operatorname{{\bf curl}}\bar{\bm{A}}$. After restriction to $\Omega$, this yields ${\bm{v}}=\operatorname{{\bf grad}}\phi$ with $\phi:=\bar{\phi}|_{\Omega}\in H^{1}(\Omega)$, respectively ${\bm{v}}=\operatorname{{\bf curl}}{\bm{A}}$ with ${\bm{A}}:=\bar{\bm{A}}|_{\Omega}\in H(\operatorname{{\bf curl}};\Omega)$. The last statement is a direct consequence of the definitions (5)–(6). ∎
2.5 Scalar potentials for the normal harmonic fields
For the space $\mathfrak{H}^{2}$ of normal harmonic fields, a natural approach is to consider scalar potentials with constant values on each connected part of the boundary. Following [31] and [1], define
$H^{1}_{S}(\Omega):=\{{\varphi}\in H^{1}(\Omega):{\varphi}|_{S_{0}}=0~\text{ and }~{\varphi}|_{S_{i}}=\text{constant},~1\leq i\leq\beta_{2}\}$ (23)
and for each $1\leq i\leq\beta_{2}$, let $\phi_{i}\in H^{1}_{S}(\Omega)$ solve the Poisson problem
$\langle\operatorname{{\bf grad}}{\varphi},\operatorname{{\bf grad}}\phi_{i}\rangle={\varphi}|_{S_{i}}\qquad\forall{\varphi}\in H^{1}_{S}(\Omega).$ (24)
This problem is clearly well-posed thanks to Poincaré’s inequality on $H^{1}_{S}(\Omega)$, which holds since $S_{0}\neq\emptyset$, and we have the following result (see, e.g. [1, Prop. 3.18]).
Proposition 2.2.
The vector fields ${\bm{v}}_{i}:=\operatorname{{\bf grad}}\phi_{i}$, $1\leq i\leq\beta_{2}$, form a basis of $\mathfrak{H}^{2}$, and it holds
$\int_{S_{j}}{\bm{n}}\cdot{\bm{v}}_{i}=\delta_{i,j},\quad 1\leq j\leq\beta_{2}.$ (25)
Proof.
We remind the argument for the sake of completeness. The first observation is that
$\operatorname{{\bf grad}}H^{1}_{S}(\Omega)\subset H_{0}(\operatorname{{\bf curl}};\Omega)$ (26)
which is easy to verify: for all $1\leq j\leq\beta_{2}$, fix a function $\psi_{j}\in H^{1}_{S}(\Omega)$ which is constant on a (disconnected) neighborhood of $\partial\Omega$ and satisfies
$\psi_{j}|_{S_{0}}=0\quad\text{ and }\quad\psi_{j}|_{S_{l}}=\delta_{j,l},\quad 1\leq l\leq\beta_{2}.$ (27)
For any ${\varphi}\in H^{1}_{S}(\Omega)$, the function $\operatorname{{\bf grad}}\big{(}{\varphi}-\sum_{j=1}^{\beta_{2}}({\varphi}|_{S_{j}})\psi_{j}\big{)}$ then belongs to $H_{0}(\operatorname{{\bf curl}};\Omega)$ as the gradient of a function in $H^{1}_{0}(\Omega)$, and its trace clearly coincides with that of $\operatorname{{\bf grad}}{\varphi}$. This essentially shows that ${\bm{v}}_{i}\in\mathfrak{H}^{2}$: indeed one obviously has $\operatorname{{\bf curl}}{\bm{v}}_{i}=0$ and $\operatorname{div}{\bm{v}}_{i}=0$ follows by taking ${\varphi}\in C^{\infty}_{c}(\Omega)$ in (24). To verify (25) we then use again the functions $\psi_{j}$ and compute
$\int_{S_{j}}{\bm{n}}\cdot{\bm{v}}_{i}=\int_{\partial\Omega}\psi_{j}~{\bm{n}}\cdot{\bm{v}}_{i}=\langle\operatorname{{\bf grad}}\psi_{j},{\bm{v}}_{i}\rangle=\langle\operatorname{{\bf grad}}\psi_{j},\operatorname{{\bf grad}}\phi_{i}\rangle=\psi_{j}|_{S_{i}}=\delta_{i,j}.$
This in particular shows that the vector fields ${\bm{v}}_{i}$ are linearly independent. The spanning property then follows from Theorem 2.1. ∎
The next section describes a construction of vector potentials for tangent harmonic fields that extends the previous ideas.
3 Construction of vector potentials for tangent harmonic fields
3.1 Tunnel curves and reciprocal surfaces
Motivated by the discussion in [8, Sec. 3], our main assumption is twofold: that the domain admits piecewise smooth “tunnel curves” $\Gamma_{i}$, $1\leq i\leq\beta_{1}$, which form a basis for the 1-chain homology group $\mathcal{H}_{1}$, and that there exist surfaces $\Sigma_{i}$ which are reciprocal in a sense specified below. An important point is that these surfaces do not need to be disjoint.
Throughout the article we will say that a function $u$ defined on a bounded Lipschitz domain $M\subset\mathbb{R}^{3}$ is piecewise $C^{k}$ if there is a covering $\overline{M}=\cup_{a=1}^{m}\overline{M}_{a}$ into disjoint open cells $M_{a}$ with Lipschitz, piecewise $C^{1}$ boundaries, such that for all $1\leq a\leq m$, the restriction $u|_{M_{a}}$ is in $C^{k}(\overline{M_{a}})$. When useful, we will denote by $C^{k}(\mathcal{M})$ the space of such functions, $\mathcal{M}=(M_{a})_{a=1}^{m}$ being the associated covering of $M$.
Assumption 3.1.
There exist closed curves $\Gamma_{i}\subset\partial\Omega$, and open oriented surfaces $\Sigma_{i}\subset\Omega$, $i=1,\ldots,\beta_{1}$, with the following properties:
  • for each $i$, there is a homeomorphism
    $T_{i}:\mathcal{O}\ni(r,\theta,z)\mapsto{\bm{x}}\in X_{i}$ (28)
    between the parametric domain $\mathcal{O}:=[0,1)\times\mathbb{T}\times(-1,1)$, where $\mathbb{T}:=\mathbb{R}\operatorname{/}\mathbb{Z}$, and some subdomain $X_{i}\subset\overline{\Omega}$, such that
    • (i)
      the surface $\{r=0\}$ parametrizes the boundary part of $X_{i}$, i.e.,
      $X_{i}\cap\partial\Omega=T_{i}(\{0\}\times\mathbb{T}\times(-1,1))$ (29)
    • (ii)
      the curve $\hat{\Gamma}:=\{r=0,z=0\}\cong\mathbb{T}$ parametrizes $\Gamma_{i}\subset X_{i}$, i.e.,
      $\Gamma_{i}=T_{i}(\{0\}\times\mathbb{T}\times\{0\})$ (30)
    • (iii)
      the mapping $T_{i}$ is piecewise $C^{\infty}$ in the above sense.
  • each surface $\Sigma_{i}$ is an open part of a smooth manifold, with a possibly disconnected boundary $\partial\Sigma_{i}$ made of one or several closed curves $\tilde{\Gamma}_{i,n}$, $n=0,\dots,\beta_{1}(\Sigma_{i})$, which are piecewise smooth and included in $\partial\Omega$.
Finally, the curves and surfaces are reciprocal in the sense that $\Gamma_{i}$ intersects $\partial\Sigma_{j}$ once if $i=j$, and zero times if $i\neq j$. More precisely, observing that $\Gamma_{i}\cap\partial\Sigma_{j}\subset X_{i}\cap\partial\Sigma_{j}$, we assume that
$X_{i}\cap\partial\Sigma_{j}=\begin{cases}T_{j}(\{0\}\times\{0\}\times(-1,1))&\text{ if }i=j,\\ \emptyset&\text{ if }i\neq j.\end{cases}$ (31)
$\mathcal{O}=[0,1)\times\mathbb{T}\times(-1,1)$
[…]
$\Omega$
$T_{i}:\mathcal{O}\to X_{i}$
[…]
Figure 2: Parametrization of a tunnel curve $\Gamma_{i}$ and its neighborhood $X_{j}$ using a piecewise smooth homeomorphism $T_{j}$ defined on a (straight) solid torus $\mathcal{O}$ periodic along $\theta$.
Remark 3.2.
The surfaces $\Sigma_{i}$ are reminiscent of the cut surfaces used to construct piecewise smooth scalar potentials in [18, 31, 1], and in the analysis they also allows us to show the linear independence of the resulting harmonic fields. One difference is that our construction does not require them to be disjoint. We note that it does not seem possible to find disjoint cut surfaces for the domain shown in Figure 1, and more generally for a hollow torus.
3.2 The surface flux functionals
Our assumptions on the surfaces $\Sigma_{i}$ guarantee that they are surface with corners [24], and more generally standard surfaces in the sense of [36, Sec. III-18], so that the Stokes theorem applies. Using the correspondence between vector fields and 1-forms, this yields
$\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{u}}$ (32)
for any vector field ${\bm{u}}\in C^{1}(\overline{\Omega})^{3}$. We observe that (32) also holds for ${\bm{u}}\in H(\operatorname{{\bf curl}};\Omega)$ which is piecewise $C^{1}$ in the sense specified above. Indeed, for every interface $f=\overline{M_{a}}\cap\overline{M_{b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, such a field satisfies
${\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{a}}={\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{b}}\quad\text{ on }f.$ (33)
Decomposing $\Sigma_{i}$ into subsurfaces $\Sigma_{i,a}=\Sigma_{i}\cap M_{a}$ with piecewise $C^{1}$ boundaries, one then applies the Stokes formula on each of these subsurfaces which yields a collection of subsurface integrals which sum to the surface integral in (32), and two types of line integrals: line integrals corresponding to interior curves (shared by two subsurfaces) $\gamma_{i,a,b}=\partial\Sigma_{i,a}\cap\partial\Sigma_{i,b}$ and line integrals over boundary curves $\gamma_{i,a}=\partial\Sigma_{i,a}\cap\partial\Sigma$. Since the former ones appear twice with opposite orientations, their sum vanishes thanks to (33), and the latter ones sum to the boundary integral in (32).
Using the same arguments as in the proof of [1, Lemma 3.10], which rely on the properties of $\Sigma_{i}$ listed in Assumption 3.1, we can extend the mapping $\gamma_{n}:C^{0}_{c}(\Omega)\to L^{\infty}(\Sigma_{i}),{\bm{w}}\mapsto{\bm{n}}\cdot{\bm{w}}|_{\Sigma_{i}}$ to a bounded trace operator $\gamma_{n}:H_{0}(\operatorname{div};\Omega)\to(H^{1/2}(\Sigma_{i}))^{\prime}.$ For each $1\leq i\leq\beta_{1}$, this allows us to consider a flux functional on $\Sigma_{i}$ defined as
$\Phi_{i}:H_{0}(\operatorname{div};\Omega)\to\mathbb{R},\quad{\bm{w}}\mapsto\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}.$ (34)
Proposition 3.3.
The linear form $\Phi_{i}$ vanishes on $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$, i.e., we have
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}})=0,\quad{\bm{v}}\in H_{0}(\operatorname{{\bf curl}};\Omega).$ (35)
Proof.
For ${\bm{v}}\in C^{\infty}_{c}(\Omega)$, Stokes’ formula (32) yields
$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}}_{n})=\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{v}}_{n}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{v}}_{n}=0$ (36)
where the last equality follows from $\partial\Sigma_{i}\subset\partial\Omega$. We conclude with a density argument: given a sequence ${\bm{v}}_{n}\in C^{\infty}_{c}(\Omega)$ converging to an arbitrary $\bm{v}$} in $H_{0}(\operatorname{{\bf curl}};\Omega)$, we have $\operatorname{{\bf curl}}{\bm{v}}_{n}\to\operatorname{{\bf curl}}{\bm{v}}$ in $H_{0}(\operatorname{div};\Omega)$ and the result follows from the continuity of the operator (34). ∎
3.3 Construction by decomposition
Our vector potentials are defined as a sum
${\bm{A}}_{i}:={\bm{A}}_{i}^{b}+{\bm{A}}_{i}^{0},\qquad 1\leq i\leq\beta_{1},$ (37)
where the first terms ${\bm{A}}_{i}^{b}$ lift some suitable boundary condition into the domain, and the second terms ${\bm{A}}_{i}^{0}$ are correction terms with homogeneous boundary conditions.
Specifically, the ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ are piecewise smooth fields (in the sense of Section 3.1) which lift the following boundary conditions:
${\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0~~\text{ on }\partial\Omega\qquad\text{ and }\qquad\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}.$ (38)
Notice that the first condition may be reformulated using a surface operator (see, e.g., [6, Cor. 3.1.16]) so that it only involves boundary values of ${\bm{A}}^{b}_{i}$ indeed. We will say that the ${\bm{A}}_{i}^{b}$ are lifted boundary potentials.
The correction potentials ${\bm{A}}_{i}^{0}$ must then satisfy
${\bm{A}}_{i}^{0}\in H_{0}(\operatorname{{\bf curl}};\Omega)\quad\text{ and }\quad\operatorname{{\bf curl}}\operatorname{{\bf curl}}{\bm{A}}_{i}^{0}=-\operatorname{{\bf curl}}\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}.$ (39)
These properties are indeed sufficient to obtain a generating family of harmonic vector potentials.
Proposition 3.4.
If (38) and (39) hold, then the potentials ${\bm{A}}_{i}$ satisfy $\operatorname{{\bf curl}}{\bm{A}}_{i}\in\mathfrak{H}^{1}$ and
$\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i})=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (40)
In particular, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$ form a basis of $\mathfrak{H}^{1}$.
Proof.
The equality $\operatorname{div}{\bm{w}}_{i}=0$ is obvious, and $\operatorname{{\bf curl}}{\bm{w}}_{i}=0$ is clear from (39). The boundary condition ${\bm{w}}_{i}\in H_{0}(\operatorname{div};\Omega)$ is also easily verified, indeed it is satisfied by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}$ thanks to (38), and by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{0}$ thanks to the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$. This shows that the fields ${\bm{w}}_{i}$ are indeed tangent-harmonic. To show the relations (40) we first consider the lifted boundary part. Since it is piecewise smooth, Stokes’ formula (32) yields
$\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i}^{b})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
where the last equality is (38). The desired relation (40) then follows from the fact that the flux functionals $\Phi_{j}$ vanish for the ${\bm{A}}^{0}_{i}$ parts, thanks to (35). The linear independence of the fields ${\bm{w}}_{i}$’s follows from (40), and the fact that they span $\mathfrak{H}^{1}$ follows from (10). ∎
We now proceed to construct the different terms of our decomposition.
3.4 Lifted boundary potentials
We construct the lifted boundary potentials by pushforward: we first define a smooth, compactly supported $\widehat{{\bm{A}}}^{b}$ on the parametric domain $\mathcal{O}$, which satisfies
$\hat{{\bm{n}}}\cdot\operatorname{{\bf curl}}\widehat{{\bm{A}}}^{b}=0\quad\text{ on }r=0,$ (41)
where $\hat{{\bm{n}}}=-\hat{{\bm{e}}}_{r}$ is the unit normal vector pointing outside of $\mathcal{O}$ on its $\{r=0\}$ boundary, and
$\int_{-1}^{1}\hat{\bm{\tau}}\cdot\widehat{{\bm{A}}}^{b}(0,0,z)\ \mathrm{d}z=1$ (42)
where $\hat{\bm{\tau}}=\hat{{\bm{e}}}_{z}$ is the unit vector tangent to the $\{r=0,\theta=0\}$ curve. Since $\widehat{{\bm{A}}}^{b}$ is compactly supported, it vanishes on every boundary of $\mathcal{O}$ with the exception of the surface $\{r=0\}$ which is contained in $\mathcal{O}$: it is on this boundary that the above boundary conditions are imposed.
Specifically, we may set
$\widehat{{\bm{A}}}^{b}(r,\theta,z):=\begin{pmatrix}0\\ 0\\ \hat{\alpha}(z)\hat{{\varphi}}(r)\end{pmatrix}$ (43)
with $\hat{\alpha}\in C_{c}^{\infty}(-1,1)$ and $\hat{{\varphi}}\in\mathcal{C}_{c}^{\infty}[0,1)$ such that $\int_{-1}^{1}\hat{\alpha}(z)\ \mathrm{d}z=1$ and $\hat{{\varphi}}(0)=1$. One easily verifies that this guarantees (41) and (42). This logical field is then pushed forward to define a piecewise smooth, compactly supported potential on the domain $X_{i}=T_{i}(\mathcal{O})$, which is then extended by zero outside of $X_{i}$. Thus, we set
${\bm{A}}_{i}^{b}({\bm{x}}):=\begin{cases}\mathcal{F}_{i}^{1}(\widehat{{\bm{A}}}^{b})({\bm{x}})~~&\text{if}~{\bm{x}}\in X_{i}\\ 0~~&\text{else}\end{cases}$ (44)
where $\mathcal{F}^{1}_{i}(\hat{\bm{u}}):=((DT_{i})^{-T}\hat{\bm{u}})\circ T_{i}^{-1}$ is the covariant Piola transform (with $DT_{i}$ the Jacobian matrix of $T_{i}$) corresponding to the pushforward operator for 1-form proxy fields [22]. Let us denote by $\mathcal{O}_{i,a}$, $a=1,\dots,m_{i}$ the subdomains of $\mathcal{O}$ on which $T_{i}$ is smooth. On the corresponding domains $X_{i,a}=T_{i}(\mathcal{O}_{i,a})$ the field ${\bm{A}}_{i}^{b}$ is smooth as well.
Proposition 3.5.
The field ${\bm{A}}_{i}^{b}$ just defined satisfies the properties stated above, namely it is $H(\operatorname{{\bf curl}};\Omega)$ and the boundary conditions (38) hold.
Proof.
To verify that ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ we need to show that (i) the tangential trace of the pushforward $\mathcal{F}_{i}^{1}(\widehat{{\bm{A}}}^{b})({\bm{x}})$ on $\partial X_{i}\setminus\partial\Omega$ vanishes, and (ii) on every interface $f=\overline{X_{i,a}}\cap\overline{X_{i,b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, we have
${\bm{\tau}}_{f}\cdot{\bm{A}}_{i}^{b}|_{X_{i,a}}={\bm{\tau}}_{f}\cdot{\bm{A}}_{i}^{b}|_{X_{i,b}}\quad\text{ on }f.$ (45)
The former property is clear since the logical field (43) vanishes on a neighborhood of $\partial\mathcal{O}\setminus\{r=0\}$, and the latter property follows from from the fact that a tangent vector on $f$ is of the form
[…]
where $\hat{\bm{\tau}}$ is a vector tangent to the logical surface $\hat{f}=T_{i}^{-1}(f)$, and where the Jacobian matrix of $T_{i}$ can be evaluated from both subdomains $\mathcal{O}_{i,a}$ or $\mathcal{O}_{i,b}$ since the directional derivative is taken along their common interface. The continuity of the tangential traces on $f$ then follows from the definition of the pushforward.
Finally to verify the boundary conditions we use the commutation relation
$\mathcal{F}^{2}_{i}\operatorname{{\bf curl}}\widehat{{\bm{A}}^{b}}=\operatorname{{\bf curl}}\mathcal{F}^{1}_{i}\widehat{{\bm{A}}^{b}}$
satisfied by the 2-form pushforward operator $\mathcal{F}^{2}_{i}$, see e.g. [22]. The relations (38) then follow from (41), (42), and the fact that line and surface integrals of smooth vector fields are respectively preserved by 1-form and 2-form pushforward operators [24, Ch. 16.6]. ∎
3.5 Correction potentials
To define the correction potentials we consider the following problem:
Find $(\sigma_{i},{\bm{A}}_{i}^{0},{\bm{p}}_{i})\in H_{0}^{1}(\Omega)\times H_{0}(\operatorname{{\bf curl}};\Omega)\times\mathfrak{H}^{2}$, such that
$\left\{\begin{aligned} \langle\sigma_{i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in H_{0}^{1}(\Omega),\\ \langle\operatorname{{\bf grad}}\sigma_{i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{i}^{b},\operatorname{{\bf curl}}{\bm{v}}\rangle,\ &&\quad\forall\ {\bm{v}}\in H_{0}(\operatorname{{\bf curl}};\Omega),\\ \langle{\bm{A}}_{i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\mathfrak{H}^{2}.\end{aligned}\right$. (46)
Proposition 3.6.
The system (46) defines a unique field ${\bm{A}}_{i}^{0}$ which satisfies (39).
Proof.
The well-posedness follows by applying [5, Theorem 3.2], to the adjoint (homogeneous) de Rham complex (2), which is possible thanks to the properties stated in [5, Sec. 4.2]. By taking respectively ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{i}$ and ${\bm{v}}={\bm{p}}_{i}\in\mathfrak{H}^{2}$, we then observe that the solution satisfies $\sigma_{i}=0$ and ${\bm{p}}_{i}=0$. The relation (39) readily follows. ∎
Combining Propositions 3.4, 3.5 and 3.6 yields the following result.
Theorem 3.1.
The fields ${\bm{A}}_{i}={\bm{A}}^{b}_{i}+{\bm{A}}^{0}_{i}$ obtained with the above construction are vector potentials for the tangent harmonic fields. In particular, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}$ that is dual to the flux functionals $\Phi_{i}$.
4 Structure-preserving discrete harmonic potentials
We end this article by describing how the above construction can be implemented in a discrete setting. The approach proposed here applies to structure-preserving finite element spaces that share a certain number of properties listed in Sections 4.1 and 4.3 below.
Assuming that the domain $\Omega$ can be triangulated with a simplicial mesh, these properties hold for instance in the case of Whitney [36, 9] and Nédélec [29] finite element spaces, and more generally for high order FEEC spaces [4], but they are not restricted to these cases.
4.1 Discretization of the homogeneous (dual) and extended dual complexes
Since harmonic fields can be described by potentials, using differential operators from the dual complex extended to spaces without homogeneous boundary conditions, a natural choice is to consider a strong discretization of the dual de Rham complex (2), seen as a subcomplex of the one without boundary conditions (i.e., the primal one).
To better reflect this choice, and distinguish this extended dual complex from the primal complex (1), let us rename the spaces with homogeneous boundary conditions as $\tilde{V}^{\ell}_{0}:=V^{*}_{3-\ell}$ with operators $\tilde{d}^{\ell}_{0}:=d^{*}_{3-\ell}$. We let then $\tilde{V}^{\ell}$ denote the corresponding spaces without boundary conditions, equipped with their natural differential operators $\tilde{d}^{\ell}$. Thus, we rewrite the dual complex (2) as
$0\xleftarrow{~~}\tilde{V}^{3}_{0}=L^{2}(\Omega)\xleftarrow{~\tilde{d}_{0}^{2}~}\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)\xleftarrow{~\tilde{d}_{0}^{1}~}\tilde{V}^{1}_{0}=H_{0}(\operatorname{{\bf curl}};\Omega)\xleftarrow{~\tilde{d}_{0}^{0}~}\tilde{V}^{0}_{0}=H^{1}_{0}(\Omega)\xleftarrow{~~}0$ (47)
and we see it as a subcomplex of
$0\xleftarrow{~~}\tilde{V}^{3}=L^{2}(\Omega)\xleftarrow{~\tilde{d}^{2}~}\tilde{V}^{2}=H(\operatorname{div};\Omega)\xleftarrow{~\tilde{d}^{1}~}\tilde{V}^{1}=H(\operatorname{{\bf curl}};\Omega)\xleftarrow{~\tilde{d}^{0}~}\tilde{V}^{0}=H^{1}(\Omega)\xleftarrow{~~}0.$ (48)
Notice that up to a change of signs in the grad and div operators, the latter coincides with the primal complex (1). For simplicity we may disregard these sign changes and consider the plain differential operators below. However, we should not confuse the primal and the extended dual complexes.
We then consider a discretization of the extended dual complex equipped with degrees of freedom that are commuting and compatible with the Dirichlet boundary conditions. By this we mean essentially two properties. First, that we are given a discrete subcomplex of (48),
$0\xleftarrow{~~}\tilde{V}^{3}_{h}\xleftarrow{~\operatorname{div}~}\tilde{V}^{2}_{h}\xleftarrow{~\operatorname{{\bf curl}}~}\tilde{V}^{1}_{h}\xleftarrow{~\operatorname{{\bf grad}}~}\tilde{V}^{0}_{h}\xleftarrow{~~}0$ (49)
with spaces of piecewise continuous functions $\tilde{V}^{\ell}_{h}\subset C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ (associated with some mesh $\mathcal{M}_{h}$ of possibly curved or polygonal Lipschitz cells, as described in Section 3.1) equipped with linear degrees of freedom
$\tilde{\sigma}^{\ell}_{i}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\mathbb{R},\qquad 1\leq i\leq N_{\ell}:=\dim(\tilde{V}^{\ell}_{h}),$ (50)
satisfying the following commuting relation: for all $v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ with $\tilde{d}^{\ell}v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell+1}$, it holds
$\tilde{\sigma}^{\ell+1}_{i}(\tilde{d}^{\ell}v)=\sum_{j=1}^{N_{\ell}}\tilde{\mathbb{D}}^{\ell}_{ij}\tilde{\sigma}^{\ell}_{j}(v),\qquad 1\leq i\leq N_{\ell+1},$ (51)
with a fixed $N_{\ell+1}\times N_{\ell}$ matrix $\mathbb{D}^{\ell}$ that represents the differential operator at the discrete level. We note that a relation of the form (51) actually implies that the corresponding projection operators $\tilde{\Pi}^{\ell}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\tilde{V}^{\ell}_{h}$, defined by
$\tilde{\sigma}^{\ell}_{i}(\tilde{\Pi}^{\ell}v)=\tilde{\sigma}^{\ell}_{i}(v),\qquad 1\leq i\leq N_{\ell},$ (52)
commute with the differential operators. The second property is that the degrees of freedom lead to projection operators that preserve the homogenous boundary conditions, in the sense that we have
$\tilde{\Pi}^{\ell}(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}_{0})=(\tilde{V}^{\ell}_{h}\cap\tilde{V}^{\ell}_{0})=:\tilde{V}^{\ell}_{h,0}.$ (53)
In particular, the same projection operators yield two commuting diagrams: one for the dual complex (with homogeneous boundary conditions)
[…]
and another one for the extended dual complex,
[…]
4.2 Flux functionals on discrete surfaces
To study the linear independence of our discrete vector potentials, we assume in addition that there exist surfaces $\Sigma_{h,i}$ made of smooth surface with corners, with the following properties: (i) the line integrals over the boundaries $\partial\Sigma_{h,i}$ can be expressed as linear combinations of degrees of freedom in $\tilde{V}^{1}_{h}$, (ii) these boundaries are included in the domain boundary ($\partial\Sigma_{h,i}\subset\partial\Omega$) and (iii) they coincide with those of the smooth reciprocal surfaces $\Sigma_{i}$ on the parametrized domains $X_{j}$, in the sense that
$X_{j}\cap\partial\Sigma_{h,i}=X_{j}\cap\partial\Sigma_{i},\qquad 1\leq i,j\leq\beta_{1}.$ (56)
We remind that these intersections are empty for $i\neq j$, see (31).
In the case of Whitney, Nédélec or more general FEEC finite elements spaces, the usual degrees of freedom involve line integrals along edges of the mesh, so that the first property holds for surfaces $\Sigma_{h,i}$ made of faces of the mesh $\mathcal{M}_{h}$. The second property poses no particular difficulty, and the last one can then be fulfilled by choosing a parametrization of $\Gamma_{i}$ such that the smooth curves $X_{i}\cap\partial\Sigma_{i}$ are included in a single edge of the boundary $\partial\Sigma_{h,i}$.
In turn, this implies that the finite element projection $\tilde{\Pi}^{1}$ interpolates the corresponding line integrals: indeed one has
$\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}(\tilde{\Pi}^{1}{\bm{a}})=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}({\bm{a}})=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$
for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$, where the coefficients $c_{i,j}$ correspond to the decomposition of the line integrals in terms of degrees of freedom. Moreover, applying Stokes’ theorem on each smooth face $f\in F(\Sigma_{h,i})$, we compute
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\sum_{f\in F(\Sigma_{h,i})}\int_{f}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{a}}_{h},\qquad{\bm{a}}_{h}\in\tilde{V}^{1}_{h},$
where the last equality uses the fact that integrals on interior curves, i.e., curves of the type $\partial f\cap\partial f^{\prime}$ with $f\neq f^{\prime}$, cancel because of their opposite orientation and the tangential continuity of the fields in $H(\operatorname{{\bf curl}};\Omega)$. As a consequence, we obtain
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=0$ (57)
for all ${\bm{a}}_{h}\in\tilde{V}^{1}_{0,h}$, and
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$ (58)
for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$.
4.3 Discrete harmonic spaces
A discretization of the harmonic spaces (5) is classically provided by the harmonic spaces associated with the discrete complex [4]. Since the primal and dual complexes have the same harmonic spaces at the continuous level, one can use the dual discrete complex (54): this leads to the discrete vector-valued spaces
$\tilde{\mathfrak{H}}^{2}_{h,0}:=\tilde{\mathfrak{Z}}^{2}_{h,0}\cap(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp},\qquad\tilde{\mathfrak{H}}^{1}_{h,0}:=\tilde{\mathfrak{Z}}^{1}_{h,0}\cap(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0})^{\perp},$ (59)
where the discrete kernel spaces are defined as usual,
$\tilde{\mathfrak{Z}}^{2}_{h,0}:=\{{\bm{w}}\in\tilde{V}^{2}_{h,0}:\operatorname{div}{\bm{w}}=0\},\qquad\tilde{\mathfrak{Z}}^{1}_{h,0}:=\{{\bm{v}}\in\tilde{V}^{1}_{h,0}:\operatorname{{\bf curl}}{\bm{v}}=0\}.$ (60)
Given the relabeling $\tilde{V}^{\ell}_{0}=V^{*}_{3-\ell}$, we note that the spaces (59) provide a discretization for the continuous harmonic spaces $\mathfrak{H}^{1}$ and $\mathfrak{H}^{2}$ respectively, so that we may also denote them as
$\mathfrak{H}^{1}_{h}:=\tilde{\mathfrak{H}}^{2}_{h,0},\qquad\mathfrak{H}^{2}_{h}:=\tilde{\mathfrak{H}}^{1}_{h,0}.$ (61)
We assume from now on that these discrete harmonic spaces have the same dimensions as their continuous counterparts, namely
$\dim(\mathfrak{H}^{1}_{h})=\beta_{1},\qquad\dim(\mathfrak{H}^{2}_{h})=\beta_{2}.$ (62)
This property holds for Whitney, Nedelec and general FEEC spaces [4, Sec. 5.5]. Finally, we can verify that discrete potentials may a priori be used for the discrete harmonic fields.
Proposition 4.1.
The following inclusions hold under the above assumptions:
$\tilde{\mathfrak{H}}^{1}_{h,0}\subseteq\big{(}\tilde{V}^{1}_{h}\cap H_{0}(\operatorname{{\bf curl}}0;\Omega)\big{)}\subseteq\operatorname{{\bf grad}}\tilde{V}^{0}_{h},$ (63)
and
$\tilde{\mathfrak{H}}^{2}_{h,0}\subseteq\big{(}\tilde{V}^{2}_{h}\cap H_{0}(\operatorname{div}0;\Omega)\big{)}\subseteq\operatorname{{\bf curl}}\tilde{V}^{1}_{h}.$ (64)
Proof.
Let ${\bm{v}}_{h}\in\tilde{\mathfrak{H}}^{1}_{h,0}$. We have ${\bm{v}}_{h}\in H_{0}(\operatorname{{\bf curl}}0;\Omega)$ by construction and the argument used to show (21) yields ${\bm{v}}_{h}=\operatorname{{\bf grad}}\phi$ for some $\phi\in H^{1}(\Omega)$. Since ${\bm{v}}_{h}$ is bounded as a piecewise continuous function, we observe that $\phi$ is also (Lipschitz) continuous, in particular we may apply the commuting projecion operators (55). This yields
${\bm{v}}_{h}=\tilde{\Pi}^{1}{\bm{v}}_{h}=\tilde{\Pi}^{1}\operatorname{{\bf grad}}\phi=\operatorname{{\bf grad}}\tilde{\Pi}^{0}\phi$
which shows (63). The embeddings (64) are proven in a similar way: For ${\bm{w}}_{h}\in\tilde{\mathfrak{H}}^{2}_{h,0}$ we have again ${\bm{w}}_{h}\in H_{0}(\operatorname{div}0;\Omega)$ by construction and the argument used to show (22) yields ${\bm{w}}_{h}=\operatorname{{\bf curl}}{\bm{A}}$ for a vector potential ${\bm{A}}\in H(\operatorname{{\bf curl}};\Omega)$ defined as the restriction of some $\bar{\bm{A}}$ on $\Omega$. Here we may consider an explicit potential defined on a cube $C_{R}=(-R,R)^{3}$ containing $\overline{\Omega}$,
$\bar{\bm{A}}({\bm{x}})=\Big{(}0,\int_{-R}^{x_{1}}\bar{w}_{h,3}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime},-\int_{-R}^{x_{1}}\bar{w}_{h,2}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime}\Big{)},$
where $\bar{\bm{w}}_{h}$ is defined as the extension of ${\bm{w}}_{h}$ by 0 on $C_{R}\setminus\Omega$. From ${\bm{w}}_{h}\in C^{0}(\mathcal{M}_{h})$ we infer that ${\bm{A}}\in C^{0}(\mathcal{M}_{h})\cap H(\operatorname{{\bf curl}};\Omega)$, so that we may again apply the commuting projection operators (55). This yields
${\bm{w}}_{h}=\tilde{\Pi}^{2}{\bm{w}}_{h}=\tilde{\Pi}^{2}\operatorname{{\bf curl}}{\bm{A}}=\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{A}}$
which proves (64). ∎
4.4 Discrete harmonic scalar potentials for $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$
To discretize the harmonic scalar potentials from Section 2.5, we can perform a simple Galerkin projection on the natural discrete subspace of $H^{1}_{S}$. Specifically, we denote
$\tilde{V}^{0}_{h,S}:=H^{1}_{S}(\Omega)\cap\tilde{V}^{0}_{h}=\{{\varphi}\in\tilde{V}^{0}_{h}:{\varphi}|_{S_{0}}=0~\text{ and }~{\varphi}|_{S_{i}}=\text{ constant},~1\leq i\leq\beta_{2}\},$ (65)
and for each $1\leq i\leq\beta_{2}$, we let $\phi_{h,i}\in\tilde{V}^{0}_{h,S}$ solve the discrete Poisson problem
$\langle\operatorname{{\bf grad}}{\varphi},\operatorname{{\bf grad}}\phi_{h,i}\rangle={\varphi}|_{S_{i}}\qquad\forall{\varphi}\in\tilde{V}^{0}_{h,S}.$ (66)
Theorem 4.1.
The fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$, $i=1,\dots,\beta_{2}$, form a basis of $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$. Moreover, they satisfy
$\langle\operatorname{{\bf grad}}\psi_{h,j},{\bm{v}}_{h,i}\rangle=\delta_{i,j},\quad 1\leq i,j\leq\beta_{2}.$ (67)
with functions $\psi_{h,j}\in\tilde{V}^{0}_{h,S}$ chosen as in Proposition 4.2 below.
Since Problem (66) is clearly well-posed, the main difficulty is to verify that the linear forms in the right-hand side are linearly independent. An answer is provided by the following result.
Proposition 4.2.
The space (65) admits a direct (a priori non orthogonal) decomposition
$\tilde{V}^{0}_{h,S}=\tilde{V}^{0}_{h,0}\oplus\bigoplus_{i=1}^{\beta_{2}}\operatorname{Span}(\psi_{h,i})$ (68)
with functions $\psi_{h,i}\in\tilde{V}^{0}_{h,S}$ that satisfy $\psi_{h,i}|_{S_{j}}=\delta_{i,j}$ for $1\leq j\leq\beta_{2}$.
Proof.
Consider the mapping $\mathcal{T}:\tilde{V}^{0}_{h,S}\to\mathbb{R}^{\beta_{2}}$, ${\varphi}\mapsto({\varphi}|_{S_{i}})_{1\leq i\leq\beta_{2}}$. Since $\ker\mathcal{T}=\tilde{V}^{0}_{h,0}$, one has
$\dim(\tilde{V}^{0}_{h,S})\leq\dim(\tilde{V}^{0}_{h,0})+\beta_{2}.$ (69)
Next, observe that any potential $\phi$ of some ${\bm{v}}\in H_{0}(\operatorname{{\bf curl}}0)$ must be constant on each surface $S_{i}$. In particular, the inclusion (63) may be refined into
$\tilde{\mathfrak{H}}^{1}_{h,0}\subseteq\operatorname{{\bf grad}}\tilde{V}^{0}_{h,S}.$ (70)
We then consider the mapping
$\mathcal{G}:\tilde{V}^{0}_{h,S}\to\tilde{\mathfrak{H}}^{1}_{h,0},\quad{\varphi}\mapsto(I-P_{\mathfrak{B}})\operatorname{{\bf grad}}{\varphi}$
where $P_{\mathfrak{B}}$ is the orthogonal projection onto $\tilde{\mathfrak{B}}^{1}_{h,0}:=\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}$. One clearly has $\tilde{V}^{0}_{h,0}\subseteq\ker\mathcal{G}$, and it follows from (70) and from the orthogonality $\tilde{\mathfrak{B}}^{1}_{h,0}\perp\tilde{\mathfrak{H}}^{1}_{h,0}$ that $\mathcal{G}$ is surjective. In particular,
$\dim(\tilde{V}^{0}_{h,S})=\dim(\ker\mathcal{G})+\dim(\tilde{\mathfrak{H}}^{1}_{h,0})\geq\dim(\tilde{V}^{0}_{h,0})+\beta_{2}$ (71)
where we have used (62) in the last step. By combining the bounds (69) and (71) one finds that $\dim(\tilde{V}^{0}_{h,S})=\dim(\tilde{V}^{0}_{h,0})+\beta_{2}$, in particular $\mathcal{T}$ is surjective which proves the desired result. ∎
Proof of Th. 4.1.
One easily verifies that the fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$ all belong to $\tilde{\mathfrak{H}}^{1}_{h,0}$ by using the same arguments as in Proposition 2.2. To see that they are linearly independent, and hence that they span $\tilde{\mathfrak{H}}^{2}_{h}$, it suffices to use ${\varphi}=\psi_{h,j}$ in (66). This also shows (67). ∎
4.5 Discrete harmonic vector potentials for $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$
The construction of discrete potentials follows the same steps as in the continuous case: we define them as a sum
${\bm{A}}_{h,i}:={\bm{A}}^{b}_{h,i}+{\bm{A}}^{0}_{h,i}\in\tilde{V}^{1}_{h},\qquad 1\leq i\leq\beta_{1},$ (72)
where the first term is a projection of the lifted boundary potential (44),
${\bm{A}}^{b}_{h,i}:=\tilde{\Pi}^{1}{\bm{A}}^{b}_{i},$ (73)
and the second term approximates the correction potential (46) by solving the corresponding discrete problem:
Find $(\sigma_{h,i},{\bm{A}}^{0}_{h,i},{\bm{p}}_{h,i})\in\tilde{V}^{0}_{h,0}\times\tilde{V}^{1}_{h,0}\times\tilde{\mathfrak{H}}^{1}_{h,0}$, such that
$\left\{\begin{aligned} \langle\sigma_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}},,\rangle\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\tilde{\mathfrak{H}}^{1}_{h,0}.\end{aligned}\right$. (74)
The well-posedness of this problem follows from the existence of uniform discrete Poincaré inequalities and bounded cochain projections, which are known to exist in the case of FEEC discretizations [4].
Observing that the flux functionals $\Phi_{i}$ from (34) are well-defined on $\tilde{\mathfrak{H}}^{2}_{h,0}$ since the latter is a subspace of $\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)$, we then have the following result.
Theorem 4.2.
The fields ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$. Moreover, their fluxes on the surfaces introduced in Section 4.2 satisfy
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot{\bm{w}}_{h,i}=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (75)
Proof.
The proof is now a matter of simple verifications. By taking ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{h,i}$ in (74) we find that $\sigma_{h,i}=0$, and with ${\bm{v}}={\bm{p}}_{h,i}$ we obtain ${\bm{p}}_{h,i}=0$. The first and last equations from (74) yield then
${\bm{A}}_{h,i}^{0}\in(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\overset{\perp}{\oplus}\tilde{\mathfrak{H}}^{1}_{h,0})^{\perp}=(\tilde{\mathfrak{Z}}^{1}_{h,0})^{\perp}$ (76)
(where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (59)–(60)), while the second equation from (74) gives
$\langle\operatorname{{\bf curl}}({\bm{A}}^{0}_{h,i}+{\bm{A}}^{b}_{h,i}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}.$ (77)
In particular, ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$ satisfies $\operatorname{div}{\bm{w}}_{h,i}=0$ and ${\bm{w}}_{h,i}\in(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp}$. As for the boundary condition ${\bm{n}}\cdot{\bm{w}}_{h,i}=0$ on $\partial\Omega$, it is verified by writing
${\bm{w}}_{h,i}=\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i}+\operatorname{{\bf curl}}{\bm{A}}^{0}_{h,i}=\tilde{\Pi}^{2}\operatorname{{\bf curl}}{\bm{A}}^{b}_{i}+\operatorname{{\bf curl}}{\bm{A}}^{0}_{h,i}$
where we have used (73) and the commutation property (55) in the second equality. Here both term satisfies the boundary condition: the first one thanks to (38) and (53), and the second one as the curl of a field in $H_{0}(\operatorname{{\bf curl}};\Omega)$. Thus we have verified that the fields ${\bm{w}}_{h,i}$ all belong to $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$ indeed.
To evaluate the fluxes across the surfaces $\Sigma_{h,j}$, we use (57) and (58): this yields
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{h,i}=\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
where in the last two steps we have used the fact that the curves $\partial\Sigma_{h,j}$ and $\partial\Sigma_{j}$ coincide on the support of ${\bm{A}}^{b}_{i}$, see (56), and the boundary conditions (38). This shows (75) and the linear independence of the fields ${\bm{w}}_{h,i}$. Their spanning properties follow from (62). ∎
4.6 Solving a simpler discrete curl-curl problem
To conclude our study and answer a question raised during a discussion with our colleague Florian Hindenlang, let us examine a simplified version of problem (74), where the discrete harmonic space $\tilde{\mathfrak{H}}^{1}_{h,0}$ is replaced by a space of the form
$\operatorname{{\bf grad}}\tilde{V}^{0}_{h,\psi}:=\operatorname{{\bf grad}}\Big{(}\bigoplus_{i=1}^{\beta_{2}}\operatorname{Span}(\psi_{h,i})\Big{)}$
associated with arbitrary functions $\psi_{h,i}\in\tilde{V}^{0}_{h}$ satisfying $\psi_{h,i}|_{S_{j}}=\delta_{i,j}$ for $0\leq j\leq\beta_{2}$. Note that the existence of such functions is established by Proposition 4.2. Thus, the new problem reads:
Find $(\sigma^{\psi}_{h,i},{\bm{A}}^{0,\psi}_{h,i},{\bm{p}}^{\psi}_{h,i})\in\tilde{V}^{0}_{h,0}\times\tilde{V}^{1}_{h,0}\times\operatorname{{\bf grad}}\tilde{V}^{0}_{h,\psi}$, such that
$\left\{\begin{aligned} \langle\sigma^{\psi}_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0,\psi}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma^{\psi}_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0,\psi},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}^{\psi}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}}\rangle,\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0,\psi},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\operatorname{{\bf grad}}\tilde{V}^{0}_{h,\psi}.\end{aligned}\right$. (78)
This problem may be of practical interest since it does not require to compute the discrete harmonic fields from $\tilde{\mathfrak{H}}^{1}_{h,0}$: the harmonic potentials from (66) may be replaced by other functions that satisfy the same piecewise constant boundary conditions but are not subject to specific constraints inside $\Omega$. In particular one may choose functions with localized supports close to the surfaces $S_{i}$. Given this freedom, there is little hope that problem (78) satisfies any stability property as the mesh is refined. Nevertheless, the following result holds.
Proposition 4.3.
Problem (78) admits a unique solution which coincides with that of problem (74), in the sense that
$(\sigma^{\psi}_{h,i},{\bm{A}}^{0,\psi}_{h,i},{\bm{p}}^{\psi}_{h,i})=(\sigma_{h,i},{\bm{A}}^{0}_{h,i},{\bm{p}}_{h,i}).$ (79)
Proof.
We will show that any solution of (78) satisfies $\sigma^{\psi}_{h,i}=0$, ${\bm{p}}^{\psi}_{h,i}=0$ and
${\bm{A}}^{0,\psi}_{h,i}\in(\tilde{\mathfrak{Z}}^{1}_{h,0})^{\perp},\qquad\langle\operatorname{{\bf curl}}({\bm{A}}_{h,i}^{0,\psi}+{\bm{A}}_{h,i}^{b}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}$ (80)
where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}=\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\oplus\tilde{\mathfrak{H}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (59)–(60). Since the relations (80) characterize a unique field in $\tilde{V}^{1}_{h,0}$, this will show the existence and uniqueness of a solution. The fact that it coincides with the solution of (74) will follow from the fact that the latter satisfies the same relations, as seen in the proof of Theorem 4.2.
So, let $\psi_{h}\in\tilde{V}^{0}_{h,\psi}$ be such that ${\bm{p}}^{\psi}_{h,i}=\operatorname{{\bf grad}}\psi_{h}$, and after observing that $\operatorname{{\bf grad}}\psi_{h}\in\tilde{V}^{1}_{h,0}$, take ${\bm{v}}=\operatorname{{\bf grad}}\sigma^{\psi}_{h,i}+{\bm{p}}^{\psi}_{h,i}=\operatorname{{\bf grad}}(\sigma^{\psi}_{h,i}+\psi_{h})$ in (78): since $\bm{v}$} is curl-free, this yields $\operatorname{{\bf grad}}(\sigma^{\psi}_{h,i}+\psi_{h})=0$, hence $\sigma^{\psi}_{h,i}+\psi_{h}=0$ given that this function must vanish on $S_{0}$. It then follows from the direct sum (68) that $\sigma^{\psi}_{h,i}=\psi_{h}=0$. In particular ${\bm{p}}^{\psi}_{h,i}=0$ and the second relation in (80) readily follows. To show the first one, observe that the first and last equation from (78) now imply
${\bm{A}}^{0,\psi}_{h,i}\in(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0})^{\perp}\cap(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,\psi})^{\perp}=(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,S})^{\perp}\subseteq(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\oplus\tilde{\mathfrak{H}}^{1}_{h,0})^{\perp}=(\tilde{\mathfrak{Z}}^{1}_{h,0})^{\perp}$
where the inclusion follows from the fact that $\tilde{\mathfrak{H}}^{1}_{h,0}$ (and $\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}$, obviously) is included in $\operatorname{{\bf grad}}\tilde{V}^{0}_{h,S}$, see (70). This shows the claimed equalities, and ends the proof. ∎
Acknowledgments
The authors would like to thank Douglas Arnold and Martin Costabel for fruitful exchanges on the de Rham theorem and its extension to Hilbert spaces. Inspiring discussions with several colleagues at IPP and outside are also acknowledged, in particular with Phil Morrison, Dirk Pauly, Patrick Ciarlet and Tobias Blickhan.

### v2

@@
 Abstract
-Representing vector fields by scalar or vector potentials can be a challenging task in domains with cavities or tunnels, due to the presence of harmonic fields which are both irrotational and solenoidal but may have no scalar or vector potentials.
-For harmonic fields normal to the boundary, which can exist in domains with cavities, it is possible to define scalar potentials with Dirichlet boundary conditions fitted to the domain’s cavities.
-For harmonic fields tangent to the boundary, which can exist in domains with tunnels, a similar construction was lacking.
-In this article we present a construction of vector potentials that yield a basis for the tangent harmonic fields.
-Our vector potentials are obtained by solving curl-curl problems with inhomogeneous tangent boundary conditions that are fitted to closed curves looping around the tunnels.
-Applied to structure-preserving finite elements, our method also provides an exact geometric parametrization of the discrete harmonic fields.
+Representing vector fields by potentials can be a challenging task in domains with cavities or tunnels, due to the presence of harmonic fields which are both irrotational and solenoidal but may have no scalar or vector potentials.
+For harmonic fields normal to the boundary, which exist in domains with cavities, the standard approach is to construct scalar potentials by solving Laplace’s equation with Dirichlet boundary conditions fitted to the closed surfaces surrounding the domain’s cavities.
+For harmonic fields tangent to the boundary, which exist in domains with tunnels, a similar method was lacking.
+In this article we present a construction of vector potentials obtained by solving curl-curl problems with inhomogeneous tangent boundary conditions fitted to closed curves looping around the tunnels.
+Just as the cavity surfaces represent a basis for the 2-chain homology group, these tunnel curves represent a basis for the 1-chain homology group and the corresponding vector potentials yield a basis for the tangent harmonic fields.
+In our analysis the linear independence of the harmonic fields is established by considering their fluxes through a collection of reciprocal surfaces.
+These surfaces, whose boundaries lie on the boundary of the domain and which are in intersection duality with the tunnel curves, represent a basis for the relative 2-chain homology group modulo the boundary: their existence in general domains follows from the Poincaré-Lefschetz duality.
+Applied to structure-preserving finite elements, our method also provides an exact geometric parametrization of the discrete harmonic fields in terms of discrete potentials.
 1 Introduction
 Using scalar or vector potentials to represent vector fields is a powerful technique in physics and engineering.
-In practical computations they ensure that a vector field is either irrotational (curl-free) or solenoidal (divergence-free), which has useful applications in many branches such as astrophysics [35, 34], magnetostatics [7, 32] or fluid dynamics [28], and they are also involved in multigrid solvers [3, 23].
+In practical computations they ensure that a vector field is either irrotational (curl-free) or solenoidal (divergence-free), which has useful applications in many branches such as astrophysics [43, 42], magnetostatics [10, 40, 32] or fluid dynamics [36], and they are also involved in multigrid solvers [5, 29].
 Beyond numerical applications, potentials are fundamental to several variational formulations, for instance in kinetic or fluid electromagnetic models [33, 35] and they play a crucial role in quantum [21] or relativistic physics [41]: a recent illustration is the elegant reformulation of the relativistic Vlasov-Maxwell equations as a wave-transport system involving Liénard-Wiechert potentials [14, 15].
-Finally, their geometric properties expressed in Stokes’ theorem relate to key topological aspects of physical laws, in particular in electromagnetism [10, 19].
+Finally, their geometric properties expressed in Stokes’ theorem are linked to key topological aspects of physical laws, particularly in electromagnetism [13, 24].
 A difficulty arises when considering domains with cavities or tunnels, where harmonic fields with vanishing traces may exist [39, 4].
 Harmonic fields are unique because they are both irrotational and solenoidal, yet they may not admit scalar or vector potentials depending on their boundary conditions.
-Specifically, domains with cavities have harmonic fields normal to the boundary that admit scalar potentials but no vector potentials, while domains with tunnels have harmonic fields tangent to the boundary that admit vector potentials but no scalar potentials.
-These cavities and tunnels correspond to the 2-chain and 1-chain homology groups of the domain, whose ranks determine the dimensions of the spaces of normal and tangent harmonic vector fields, respectively [24, 2]: in particular these dimensions are topological invariants, which means that they remain unchanged under continuous deformations (with continuous inverse) of the domain.
-For normal harmonic fields, a construction has been proposed for scalar potentials, characterized by Laplace equations with boundary conditions fitted to the domain’s cavities [31].
-For tangent harmonic fields however, a similar construction is more complex: in domains diffeomorphic to a solid torus, a common approach is to use the toroidal angle as a (piecewise smooth) scalar potential [26, 20], and this approach has been extended to a large class of domains, namely domains that can be made simply connected by removing a finite number of disjoint cuts surfaces [18, 31, 1].
-A construction of proper vector potentials, however, seems to be lacking for tangent harmonic fields.
+Specifically, domains with cavities have harmonic fields normal to the boundary that admit scalar potentials but no vector potentials, while domains with tunnels have harmonic fields tangent to the boundary that admit vector potentials but no scalar potentials [16].
+These cavities and tunnels correspond to the 2-chain and 1-chain homology groups of the domain, whose ranks (the Betti numbers) coincide with the dimensions of the spaces of normal and tangent harmonic vector fields, respectively [30, 4]: in particular these dimensions are topological invariants, which means that they remain unchanged under homeomorphisms.
+For normal harmonic fields, a canonical construction has been proposed where scalar potentials are characterized by Laplace problems with boundary conditions fitted to the domain’s cavities [39].
+For tangent harmonic fields, a similar construction is more complex: in domains diffeomorphic to a solid torus, a common approach is to use the toroidal angle as a (piecewise smooth) scalar potential [34, 26], and this approach has been extended to domains that can be made simply connected by removing a finite number of disjoint cutting surfaces [22, 39, 3].
+Such an assumption, however, imposes topological constraints which exclude simple domains such as hollow tori [9].
+Moreover, a construction of proper vector potentials seemed to be lacking for tangent harmonic fields.
 In this article we propose such a construction.
-Our vector potentials are characterized by homogeneous curl-curl problems, with specific tangent boundary conditions fitted to closed curves that loop around the tunnels of the domain.
-These “tunnel curves” form a generating family for the 1-chain homology group, and the associated vector potentials yield a basis for the space of tangent harmonic vector fields.
-In our analysis we assume the existence of surfaces reciprocal to the tunnel curves: these surfaces are reminiscent of the cut surfaces involved in the construction of piecewise smooth scalar potentials, but here their purpose is only to establish the linear independence of the resulting harmonic fields.
-In particular they do not need not be disjoint, which allows us to apply our construction to domains such as hollow tori.
-An interesting feature of our approach is that it can also be applied in a structure-preserving finite element framework.
-Under some assumptions which are fulfilled by Whitney [36, 9] and higher order Nédélec elements [29, 30], and more generally by the discrete de Rham sequences constructed by Hiptmair [21] and Arnold, Falk and Winther [4, 5] in the Finite Element Exterior Calculus (FEEC) framework, a natural discretization of our construction leads to an exact geometric parametrization of the discrete harmonic fields.
+Our vector potentials are characterized by homogeneous curl-curl problems with tangent boundary conditions fitted to closed curves that loop around the tunnels of the domain.
+These tunnel curves form a basis family for the 1-chain homology group, and the associated vector potentials yield a basis for the space of tangent harmonic vector fields.
+To establish the linear independence of the resulting harmonic fields, our analysis relies on surfaces which are reciprocal to the tunnel curves.
+These surfaces are reminiscent of the cutting surfaces involved in the construction of piecewise smooth scalar potentials, however they are not involved in the construction step and for this reason they do not need to be disjoint or render the complementary domain simply connected.
+Instead, their main properties are that their own boundaries lie on the boundaries of the domain, and that they realize an intersection duality pairing with the tunnel curves.
+The fact that such a construction is possible for general domains is a consequence of the classical Poincaré-Lefschetz duality [11, 23].
+At the discrete level, we note that several authors have developed tree-based approaches as an alternative to cutting surfaces, which have been used to compute bases of divergence-free FEM spaces in [25, 2], vector potentials of (non harmonic) prescribed vector fields in [40], and discrete loop fields in [1], which may be used to generate a basis for tangent harmonic fields.
+An interesting feature of our approach is that it can also be applied to compute discrete harmonic vector potentials in structure-preserving finite element spaces.
+Under some assumptions which are fulfilled by Whitney [44, 12] or higher order Nédélec elements [37, 38], and more generally by the discrete de Rham sequences constructed by Hiptmair [27] and Arnold, Falk and Winther [6, 7] in the Finite Element Exterior Calculus (FEEC) framework, a natural discretization of our construction leads to an exact geometric parametrization of the discrete harmonic fields.
 The article is organized as follows.
 In Section 2 we begin by reminding the definition of the homology groups, together with that of the harmonic spaces in the $L^{2}$ de Rham complex, and we discuss some important properties regarding their dimensions and potentials.
-In Section 3 we then describe our construction, which is based on a decomposition of the vector potentials in two parts: a smooth part that lifts a specific boundary condition fitted to the tunnel curves, and a correction part that guarantees that the resulting potential solves an homogeneous curl-curl problem.
+In Section 3 we then discuss our main assumption and describe our construction, which is based on a decomposition of the vector potentials in two parts: a smooth part that lifts a specific boundary condition fitted to the tunnel curves, and a correction part that guarantees that the resulting potential solves an homogeneous curl-curl problem.
 Section 4 is finally devoted to studying a discrete version of our construction.
@@
 For instance, a construction of discrete vector potentials for magnetic fields satisfying specific flux conditions is described in [13], within the framework of Whitney elements.
-We learned of Alain Bossavit’s passing as we were completing this work, which we dedicate to his memory.
+We learned of Alain Bossavit’s passing as we were writing this article, which we dedicate to his memory.
 2 Homology groups and harmonic function spaces
-Throughout the article, we consider that $\Omega$ is an open bounded domain in $\mathbb{R}^{3}$ with a Lipschitz boundary, in the usual sense that $\partial\Omega$ is locally the graph of Lipschitz mappings, with $\Omega$ on one side only.
+Throughout the article, $\Omega$ is an open bounded domain in $\mathbb{R}^{3}$ with Lipschitz boundary, in the usual sense that $\partial\Omega$ is locally the graph of a Lipschitz mapping with $\Omega$ on one side only.
 For simplicity, we assume that $\Omega$ is connected.
@@
 which yields for the Hilbert de Rham complex above:
-$\left\{\begin{aligned} &\mathfrak{H}^{0}=\mathbb{R},\\ &\mathfrak{H}^{1}=H(\operatorname{{\bf curl}}0;\Omega)\cap H_{0}(\operatorname{div}0;\Omega),\\ &\mathfrak{H}^{2}=H(\operatorname{div}0;\Omega)\cap H_{0}(\operatorname{{\bf curl}}0;\Omega),\\ &\mathfrak{H}^{3}=\{0\}\end{aligned}\right$.
+$\left\{\begin{aligned} &\mathfrak{H}^{0}=\mathbb{R},\\ &\mathfrak{H}^{1}=H(\operatorname{{\bf curl}}0;\Omega)\cap H_{0}(\operatorname{div}0;\Omega),\\ &\mathfrak{H}^{2}=H(\operatorname{div}0;\Omega)\cap H_{0}(\operatorname{{\bf curl}}0;\Omega),\\ &\mathfrak{H}^{3}=\{0\}.\end{aligned}\right$.
 (6)
-where we have used standard concise notation for the kernel spaces of the curl and divergence operators with or without boundary conditions, see e.g.
+Here, we have used standard concise notation for the kernel spaces of the curl and divergence operators with or without boundary conditions, see e.g.
 [8].
+In (6) we see two spaces of harmonic vector fields: $\mathfrak{H}^{1}$ consists of fields which are tangent to the boundary (as their normal trace vanishes), whereas $\mathfrak{H}^{2}$ consists of fields which are normal to it.
 If the domain is contractible (i.e., simply connected with a connected boundary), the above spaces of harmonic vector fields must vanish as a consequence of the Poincaré lemma [13].
@@
 The works of de Rham [20] have established profound connections between harmonic spaces and homology groups.
-Here we will follow [24] and consider singular homology groups.
-We remind that a singular $k$-simplex is a continuous map $\sigma:\Delta_{k}\to\Omega$ defined on the standard $k$-simplex $\Delta_{k}$, and that a (singular) $k$-chain is a formal linear combination of such singular simplices, with coefficients in $\mathbb{Z}$.
-Denoting by $\mathcal{C}_{k}(\Omega)$ the corresponding group of $k$-chains, and by $\partial:\mathcal{C}_{k}(\Omega)\to\mathcal{C}_{k-1}(\Omega)$ the (singular) boundary operator, we obtain a chain complex by considering the sequence
-$0\xleftarrow{~~}\mathcal{C}_{0}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{1}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{2}(\Omega)\xleftarrow{~\partial~}\mathcal{C}_{3}(\Omega)\xleftarrow{~~}0.$ (7)
+Here we follow [23] and consider singular homology groups in the compact domain $\overline{\Omega}$ (a topological manifold).
+We remind that a singular $k$-simplex (with $0\leq k\leq 3$ here) is a continuous map $\sigma:\Delta_{k}\to\overline{\Omega}$ defined on the standard $k$-simplex $\Delta_{k}$, and that a (singular) $k$-chain is a formal linear combination of such singular simplices, with coefficients in $\mathbb{Z}$.
+Denoting by $\mathcal{C}_{k}(\overline{\Omega})$ the corresponding group of $k$-chains, and by $\partial:\mathcal{C}_{k}(\overline{\Omega})\to\mathcal{C}_{k-1}(\overline{\Omega})$ the (singular) boundary operator, we obtain a chain complex by considering the sequence
+$0\xleftarrow{~~}\mathcal{C}_{0}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{1}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{2}(\overline{\Omega})\xleftarrow{~\partial~}\mathcal{C}_{3}(\overline{\Omega})\xleftarrow{~~}0.$ (7)
 Here, the order from right to left corresponds to the orientation of the dual complex (2).
-Denoting by $\mathcal{Z}_{k}$ and $\mathcal{B}_{k}$ the $k$-cycles and $k$-boundaries, which are the kernels and ranges (in $\mathcal{C}_{k}(\Omega)$) of the boundary operator, and observing that the inclusion $\mathcal{B}_{k}\subseteq\mathcal{Z}_{k}$ follows from the sequence property $\partial\circ\partial=0$, one next defines the $k$-th homology group as the quotient
-$\mathcal{H}_{k}:=\mathcal{Z}_{k}\operatorname{/}\mathcal{B}_{k}.$ (8)
-Thus, an element of $\mathcal{H}_{k}$ is a class of closed $k$-chains (closed in a manifold sense, not a topological one) modulo boundary chains.
-Intuitively, this corresponds to the set of points, closed curves or closed surfaces (there are no closed volumes in $\Omega$) that can be obtained from each other by continuous deformations within the domain $\Omega$.
+Denoting by $\mathcal{Z}_{k}(\overline{\Omega})$ and $\mathcal{B}_{k}(\overline{\Omega})$ the $k$-cycles and $k$-boundaries, which are the kernels and ranges (in $\mathcal{C}_{k}(\overline{\Omega})$) of the boundary operator, and observing that the inclusion $\mathcal{B}_{k}(\overline{\Omega})\subseteq\mathcal{Z}_{k}(\overline{\Omega})$ follows from the sequence property $\partial\circ\partial=0$, one next defines the $k$-th homology group as the quotient
+$\mathcal{H}_{k}(\overline{\Omega}):=\mathcal{Z}_{k}(\overline{\Omega})\operatorname{/}\mathcal{B}_{k}(\overline{\Omega}).$ (8)
+Thus, an element of $\mathcal{H}_{k}(\overline{\Omega})$ is a class $\sigma]=\{\sigma+\partial\tau:\tau\in\mathcal{C}_{k+1}(\overline{\Omega})\}$ of closed $k$-chains $\sigma$ (closed in a manifold sense, not a topological one), i.e.
+cycles, modulo boundary chains.
+Intuitively, this corresponds to the set of points (closed by convention), closed curves or closed surfaces that can be obtained from each other by continuous deformations within $\overline{\Omega}$.
+Note that in $\mathbb{R}^{3}$ no bounded volume can be closed in a manifold sense.
 An illustration is provided in Figure 1.
@@
 The domain boundary has three connected components which are closed surfaces (in orange, only partially drawn for the sake of visibility): $S_{0}$ is the “outer” boundary which bounds the unbounded component of $\mathbb{R}^{3}\setminus\Omega$, $S_{1}$ bounds the contractible cavity, and $S_{2}$ bounds the inner tunnel.
-The last two represent independent classes of the 2-chain homology group $\mathcal{H}_{2}$, while the first one is homologous to $(S_{1}+S_{2})$.
-Finally, the closed curves $\Gamma_{i}\subset\partial\Omega$, $i=1,2$, (in light blue) represent two independent classes of the 1-chain homology group $\mathcal{H}_{1}$.
+The last two represent independent classes of the 2-chain homology group $\mathcal{H}_{2}(\overline{\Omega})$, while the first one is homologous to $(S_{1}+S_{2})$.
+Finally, the closed curves $\Gamma_{i}\subset\partial\Omega$, $i=1,2$, (in light blue) represent two independent classes of the 1-chain homology group $\mathcal{H}_{1}(\overline{\Omega})$.
 They admit reciprocal surfaces $\Sigma_{i}$ as described in Assumption 3.3, visualized by their boundary curves $\tilde{\Gamma}_{i,a}\subset\partial\Omega$, drawn in red.
-The rank of the $k$-th homology group is called the $k$-th Betti number of $\Omega$, and denoted
-$\beta_{k}:=\dim(\mathcal{H}_{k}).$ (9)
+One can verify that $\mathcal{H}_{k}(\overline{\Omega})$ is a free abelian group, akin to a vector space with integer coefficients. Its rank is called the $k$-th Betti number of $\Omega$, and denoted
+$\beta_{k}:=\operatorname{rank}(\mathcal{H}_{k}(\overline{\Omega})).$ (9)
+Remark 2.1.
+One may also consider singular chains in the open domain $\Omega$, and the corresponding homology groups $\mathcal{H}_{k}(\Omega)$.
+By considering a small perturbation that moves every point of $\overline{\Omega}$ towards the interior of $\Omega$, it is not difficult to see that the two homology groups have the same ranks.
 The extremal numbers are simple: we always have $\beta_{3}=0$ since there are no closed volumes (three-dimensional cycles) in $\Omega$, and $\beta_{0}$ is the number of connected components, hence $\beta_{0}=1$ for a connected domain.
@@
 In particular, $\beta_{2}$ is the number of cavities in the domain, i.e., the number of connected components of $\mathbb{R}^{3}\setminus\Omega$ which are bounded, and $\beta_{1}$ is the number of tunnels.
-Notice that tunnels may go through $\Omega$ like handles, or be entirely surrounded by it, which makes them cavities as well: an example is shown in Figure 1.
-In the sequel we will denote by $S_{i}$, $i=0,\dots,\beta_{2}$, the connected components of the boundary $\partial\Omega$ (oriented by the normal vector pointing outside of $\Omega$).
-They are all closed surfaces: we let $S_{0}$ be the boundary of the single unbounded connected component of $\mathbb{R}^{3}\setminus\Omega$, so that for every $i=1,\dots,\beta_{2}$, $S_{i}$ bounds a single cavity surrounded by the domain $\Omega$: these surfaces form a basis of the 2-chain homology group $\mathcal{H}_{2}$.
+Notice that tunnels may go through $\Omega$ as in handles, or be entirely surrounded by it, which makes them cavities as well: an example of both types is shown in Figure 1.
+In the sequel we will denote by $S_{i}$, $i=0,\dots\beta_{2}$, the connected components of the boundary $\partial\Omega$ (oriented by the normal vector pointing outside of $\Omega$).
+They are all closed surfaces: we let $S_{0}$ be the boundary of the single unbounded connected component of $\mathbb{R}^{3}\setminus\Omega$, so that for every $i=1,\dots\beta_{2}$, $S_{i}$ bounds a single cavity surrounded by the domain $\Omega$: these surfaces form a basis of the 2-chain homology group $\mathcal{H}_{2}$.
 Observe that $S_{0}$ is homologous to $(\sum_{i=1}^{\beta_{2}}S_{i})$ since the difference bounds the domain itself.
-Similarly, we will denote by $\Gamma_{i}$, $i=1,\dots,\beta_{1}$, a collection of oriented closed curves that generate the 1-chain homology group $\mathcal{H}_{1}$.
+Similarly, we will denote by $\Gamma_{i}$, $i=1,\dots\beta_{1}$, a family of oriented closed curves that generate the 1-chain homology group $\mathcal{H}_{1}$.
 These curves will be chosen on the boundary $\partial\Omega$, such that each one circles once around a different tunnel: a precise description will be discussed in Section 3.1 below.
@@
 Theorem 2.1.
-For $k=0,\dots,3$, the following relation holds:
+For $k=0,\dots 3$, the following relation holds:
 $\dim(\mathfrak{H}^{k})=\beta_{k}.$ (10)
-Since the classical version of de Rham’s theorem is formulated for the smooth complex, we give here a complete argument which uses a result of Costabel and McIntosh [14], suggested to us by Douglas Arnold.
+Since the classical version of de Rham’s theorem is formulated for the smooth complex, we give here a detailed argument which uses a result of Costabel and McIntosh [18], suggested to us by Douglas Arnold.
 Proof.
-As an open subset of $\mathbb{R}^{3}$, $\Omega$ is a smooth manifold and the de Rham theorem applies, for instance as formulated in [24, Th.
-18.14].
-As a consequence, the $k$-chain cohomology groups (isomorphic to the dual spaces $(\mathcal{H}_{k})^{\prime}$ and hence of the same dimensions $\beta_{k}$) are isomorphic to the cohomology groups of the smooth de Rham complex
+According to Remark 2.1 we may consider the homology groups on $\Omega$ which, as an open subset of $\mathbb{R}^{3}$, is a smooth manifold (without boundary).
+We can then apply the de Rham theorem as formulated in [30, Th.
+18.14], which states that the $k$-chain cohomology groups $\mathcal{H}^{k}(\Omega;\mathbb{R})$ (which are real vector spaces isomorphic to the dual spaces $(\mathcal{H}_{k}(\Omega))^{\prime}$ of linear forms on $\mathcal{H}_{k}(\Omega)$, and hence of the same dimensions $\beta_{k}$) are isomorphic to the cohomology groups (spaces) of the smooth de Rham complex
 $0\xrightarrow{~~}C^{\infty}(\Omega)\xrightarrow{~\operatorname{{\bf grad}}~}C^{\infty}(\Omega)^{3}\xrightarrow{~\operatorname{{\bf curl}}~}C^{\infty}(\Omega)^{3}\xrightarrow{~\operatorname{div}~}C^{\infty}(\Omega)\xrightarrow{~~}0.$ (11)
@@
 $\dim\big(\mathfrak{Z}^{k}_{\infty}\operatorname{/}\mathfrak{B}^{k}_{\infty}\big)=\beta_{k}.$ (12)
-Moreover, a Poincaré duality theorem holds ([24, Pbm.
+Moreover, a Poincaré duality theorem holds ([30, Pbm.
 18.7]), which states that the cohomology groups of the compactly supported de Rham sequence (indexed with reverse order) have the same dimensions.
@@
 With regard to potentials, the question is then to know whether harmonic vector fields belong to some range space.
-It is clear from (19) that the tangential harmonic fields in $\mathfrak{H}^{1}$ have no scalar potential, from (20) we see that the normal harmonic fields in $\mathfrak{H}^{2}$ have no vector potential.
+It is clear from (19) that the tangential harmonic fields in $\mathfrak{H}^{1}$ have no scalar potential, and from (20) we see that the normal harmonic fields in $\mathfrak{H}^{2}$ have no vector potential.
 But the reverse may be true, and a simple argument shows that it is indeed.
@@
 ∎
+Remark 2.3.
+For a characterization of smooth fields with scalar or vector potentials, see [16].
 2.5 Scalar potentials for the normal harmonic fields
@@
 $\langle\operatorname{{\bf grad}}{\varphi},\operatorname{{\bf grad}}\phi_{i}\rangle={\varphi}|_{S_{i}}\qquad\forall{\varphi}\in H^{1}_{S}(\Omega).$ (24)
-This problem is clearly well-posed thanks to Poincaré’s inequality on $H^{1}_{S}(\Omega)$, which holds since $S_{0}\neq\emptyset$, and we have the following result (see, e.g. [1, Prop. 3.18]).
+This problem is clearly well-posed thanks to Poincaré’s inequality on $H^{1}_{S}(\Omega)$, which holds since $S_{0}\neq\emptyset$, and we have the following result (see, e.g. [3, Prop. 3.18]).
 Proposition 2.4.
@@
 $\operatorname{{\bf grad}}H^{1}_{S}(\Omega)\subset H_{0}(\operatorname{{\bf curl}};\Omega)$ (26)
-which is easy to verify: for all $1\leq j\leq\beta_{2}$, fix a function $\psi_{j}\in H^{1}_{S}(\Omega)$ which is constant on a (disconnected) neighborhood of $\partial\Omega$ and satisfies
+which is easy to verify: for all $1\leq j\leq\beta_{2}$, fix a function $\psi_{j}\in H^{1}_{S}(\Omega)$ that is constant on a (disconnected) neighborhood of $\partial\Omega$ and satisfies
 $\psi_{j}|_{S_{0}}=0\quad\text{ and }\quad\psi_{j}|_{S_{l}}=\delta_{j,l},\quad 1\leq l\leq\beta_{2}.$ (27)
@@
 3.1 Tunnel curves and reciprocal surfaces
-Motivated by the discussion in [8, Sec.
-3], our main assumption is twofold: that the domain admits piecewise smooth “tunnel curves” $\Gamma_{i}$, $1\leq i\leq\beta_{1}$, which form a basis for the 1-chain homology group $\mathcal{H}_{1}$, and that there exist surfaces $\Sigma_{i}$ which are reciprocal in a sense specified below.
-An important point is that these surfaces do not need to be disjoint.
-Throughout the article we will say that a function $u$ defined on a bounded Lipschitz domain $M\subset\mathbb{R}^{3}$ is piecewise $C^{k}$ if there is a covering $\overline{M}=\cup_{a=1}^{m}\overline{M}_{a}$ into disjoint open cells $M_{a}$ with Lipschitz, piecewise $C^{1}$ boundaries, such that for all $1\leq a\leq m$, the restriction $u|_{M_{a}}$ is in $C^{k}(\overline{M_{a}})$.
-When useful, we will denote by $C^{k}(\mathcal{M})$ the space of such functions, $\mathcal{M}=(M_{a})_{a=1}^{m}$ being the associated covering of $M$.
+Motivated by the discussion in [11, Sec.
+3], our construction relies on the assumption that the domain $\Omega$ admits piecewise smooth “tunnel curves” $\Gamma_{i}$, which are closed curves lying on the boundary $\partial\Omega$, and “reciprocal surfaces” $\Sigma_{i}\subset\overline{\Omega}$, $1\leq i\leq\beta_{1}$, whose boundaries $\partial\Sigma_{i}$ also lie on the domain boundary and intersect the tunnel curves in a dual manner, in the sense that the number of intersection points of $\Gamma_{i}$ and $\Sigma_{j}$ satisfies
+$I(\Gamma_{i},\Sigma_{j})=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (28)
+Aside from regularity considerations, the fact that such a property holds for general domains follows from classical results in algebraic topology, and in particular from the geometric realization of Poincaré-Lefschetz duality through intersection pairings.
+This important duality result states that the $k$-chain homology groups $\mathcal{H}_{k}(\overline{\Omega})$ are isomorphic to the (3k)(3-k)-chain relative homology groups $\mathcal{H}_{3-k}(\overline{\Omega},\partial\Omega)$, which are the homology groups of the singular chains modulo the boundary – two chains being equal modulo the boundary if their difference lies on $\partial\Omega$, see e.g.
+[31, 11].
+A modern exposition of these results can be found in the detailed lecture notes of Stefan Friedl [23]: applying Theorem 194.14 and Proposition 198.2 one finds indeed that there exists a non-singular pairing between the above (torsion-free) homology groups,
+$Q=Q^{as}_{\overline{\Omega},\emptyset,\partial\Omega}:\mathcal{H}_{k}(\overline{\Omega})\times\mathcal{H}_{3-k}(\overline{\Omega},\partial\Omega)\to\mathbb{Z}~,$ (29)
+called asymmetric intersection pairing for the reason that for any pair of cycles $\gamma\in\mathcal{Z}_{k}(\overline{\Omega})$ and $\sigma\in\mathcal{Z}_{3-k}(\overline{\Omega},\partial\Omega)$ which intersect nicely, $Q([\gamma],[\sigma])$ coincides with the algebraic intersection number of $\gamma$ and $\sigma$.
+With $k=1$, we thus find that for any collection of (reasonably smooth) closed curves $\Gamma_{i}$ representing a basis of $\mathcal{H}_{1}(\overline{\Omega})$, there exists a collection of surfaces $\Sigma_{i}$ which are closed modulo to the boundary, represent a basis of $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$ and are such that $Q([\Gamma_{i}],[\Sigma_{j}])=\delta_{i,j}$ for all $i,j$.
+We then observe that the curves $\Gamma_{i}$ may be chosen to lie on the domain boundary without changing their homology class $\Gamma_{i}]$, and that the surfaces fulfill the above requirements (as long as they can be chosen to intersect nicely the curves, which is not a restrictive assumption in a Lipschitz domain $\Omega$): being closed modulo the boundary means indeed that $\partial\Sigma_{i}\subset\partial\Omega$, and a relation of the form (28) follows directly from the $Q$-duality of the homology classes, with the specification that $I$ denotes the algebraic intersection number of the singular chains, which weighs each intersection with a sign corresponding to their respective orientations.
+$\mathcal{O}=[0,1)\times\mathbb{T}\times(-1,1)$
+[…]
+[…]
+$\Omega$
+$T_{i}:\mathcal{O}\to X_{i}$
+[…]
+[…]
+[…]
+Figure 2: Illustration of a tunnel curve $\Gamma_{i}$ and its neighborhood $X_{i}$, parametrized with a piecewise smooth homeomorphism $T_{i}$ defined on the (straight) torus $\mathcal{O}$, periodic along the variable $\theta$.
+To formulate our precise working hypotheses, we now introduce a couple of definitions.
+Definition 3.1.
+A function $u$ defined on a bounded Lipschitz domain $M\subset\mathbb{R}^{d}$ is said piecewise $C^{k}$, $k\geq 0$, if there exists a decomposition $\overline{M}=\cup_{a=1}^{m}\overline{M}_{a}$ into disjoint open cells $M_{a}$ with Lipschitz, piecewise $C^{1}$ boundaries, such that for all $1\leq a\leq m$, the restriction $u|_{M_{a}}$ is in $C^{k}(\overline{M_{a}})$.
+If in addition $u$ is invertible and $u^{-1}$ is piecewise $C^{k}$ over the cells $u(M_{a})$, $1\leq a\leq m$, then we say that $u$ is a piecewise $C^{k}$ diffeomorphism.
+In the sequel we will refer to such a decomposition $\mathcal{M}=\{M_{a}:a=1,\dots m\}$ as a mesh, and we will denote by $C^{k}(\mathcal{M})$ the space of piecewise $C^{k}$ functions over $\mathcal{M}$.
+Definition 3.2.
+We say that a surface $\Sigma$ and a mesh $\mathcal{M}$ are compatible if there exists a decomposition $\Sigma=\cup_{a=1}^{n}\Sigma_{a}$ made of smooth surfaces with corners [30], such that the interiors $\Sigma_{a}\setminus\partial\Sigma_{a}$ are disjoint and included in closed mesh cells $\overline{M}_{b}$ for some $b=b(a)\in\{0,\dots m\}$.
+Our working hypotheses read then as follows.
 Assumption 3.3.
-There exist closed curves $\Gamma_{i}\subset\partial\Omega$, and open oriented surfaces $\Sigma_{i}\subset\Omega$, $i=1,\ldots,\beta_{1}$, with the following properties:
-  • for each $i$, there is a homeomorphism
+There exist oriented curves $\Gamma_{i}\subset\partial\Omega$ and surfaces $\Sigma_{i}\subset\overline{\Omega}$, $i=1,\ldots\beta_{1}$, with the following properties:
+  • Each curve $\Gamma_{i}$ is closed and piecewise smooth, in particular there is a homeomorphism
     $T_{i}:\mathcal{O}\ni(r,\theta,z)\mapsto{\bm{x}}\in X_{i}$ (30)
@@
       $\Gamma_{i}=T_{i}(\{0\}\times\mathbb{T}\times\{0\})$ (32)
+      and $\Gamma_{i}$ is oriented by the unit tangent vector ${\bm{\tau}}_{\Gamma_{i}}$ along increasing values of $\theta$
     • (iii)
-      the mapping $T_{i}$ is piecewise $C^{\infty}$ in the above sense.
-  • each surface $\Sigma_{i}$ is an open part of a smooth manifold, with a possibly disconnected boundary $\partial\Sigma_{i}$ made of one or several closed curves $\tilde{\Gamma}_{i,n}$, $n=0,\dots,\beta_{1}(\Sigma_{i})$, which are piecewise smooth and included in $\partial\Omega$.
-Finally, the curves and surfaces are reciprocal in the sense that $\Gamma_{i}$ intersects $\partial\Sigma_{j}$ once if $i=j$, and zero times if $i\neq j$.
-More precisely, observing that $\Gamma_{i}\cap\partial\Sigma_{j}\subset X_{i}\cap\partial\Sigma_{j}$, we assume that
-$X_{i}\cap\partial\Sigma_{j}=\begin{cases}T_{j}(\{0\}\times\{0\}\times(-1,1))&\text{ if }i=j,\\ \emptyset&\text{ if }i\neq j.\end{cases}$ (31)
-$\mathcal{O}=[0,1)\times\mathbb{T}\times(-1,1)$
-[…]
-[…]
-$\Omega$
-$T_{i}:\mathcal{O}\to X_{i}$
-[…]
-[…]
-[…]
-Figure 2: Parametrization of a tunnel curve $\Gamma_{i}$ and its neighborhood $X_{j}$ using a piecewise smooth homeomorphism $T_{j}$ defined on a (straight) solid torus $\mathcal{O}$ periodic along $\theta$.
+      the mapping $T_{i}$ is a piecewise $C^{\infty}$ diffeomorphism in the sense of Definition 3.1.
+Letting $\mathcal{O}_{i,a}$, $a=1,\dots m_{i}$, be the associated open mesh cells of $\mathcal{O}$, we denote by
+      $\mathcal{X}_{i}=\{X_{i,a}:a=0,\dots m_{i}\}\quad\text{with}\quad\begin{cases}X_{i,0}:=\Omega\setminus\overline{X}_{i},\\ X_{i,a}:=T_{i}(\mathcal{O}_{i,a}),~~a=1,\dots m_{i},\end{cases}$ (33)
+      the corresponding mesh of $\Omega$.
+  • Each surface $\Sigma_{i}$ is Lipschitz, compatible with all the meshes $\mathcal{X}_{j}$, $j=1,\dots\beta_{1}$, in the sense of Definition 3.2, and its (manifold) boundary $\partial\Sigma_{i}$ is included in $\partial\Omega$.
+Finally, the above curves and surfaces are reciprocal in the sense that $\Gamma_{i}$ intersects $\Sigma_{j}$ once if $i=j$, and zero times if $i\neq j$.
+More precisely, observing that $\Gamma_{i}\cap\Sigma_{j}=\Gamma_{i}\cap\partial\Sigma_{j}\subset X_{i}\cap\partial\Sigma_{j}$, it holds
+$X_{i}\cap\partial\Sigma_{j}=\begin{cases}T_{i}\big(\{0\}\times\{0\}\times(-1,1)\big)&\text{ if }i=j,\\ \emptyset&\text{ if }i\neq j,\end{cases}$ (34)
+and we orient each surface $\Sigma_{i}$ by letting its unit normal vector ${\bm{n}}_{\Sigma_{i}}$ at the intersection point $P_{i}=\Gamma_{i}\cap\Sigma_{i}=T_{i}(0,0,0)$ satisfy ${\bm{n}}_{\Sigma_{i}}\cdot{\bm{\tau}}_{\Gamma_{i}}>0$.
 Remark 3.4.
-The surfaces $\Sigma_{i}$ are reminiscent of the cut surfaces used to construct piecewise smooth scalar potentials in [18, 31, 1], and in the analysis they also allows us to show the linear independence of the resulting harmonic fields.
-One difference is that our construction does not require them to be disjoint.
-We note that it does not seem possible to find disjoint cut surfaces for the domain shown in Figure 1, and more generally for a hollow torus.
+In the above assumption, and in particular in (34), we have considered for simplicity that an intersection duality of the form (28) can be realized with geometric intersection numbers $I(\Gamma_{i},\Sigma_{j})=\#(\Gamma_{i}\cap\Sigma_{j})$, which is the case for virtually any domain we can think of.
+In the general case however, this duality is only established with algebraic intersection numbers as discussed above.
+Instead of (34), one should thus assume that for any $i,j$, there exists distinct points $\theta(i,j,p)\in\mathbb{T}$, $1\leq p\leq n(i,j)$ (with $n(i,j)$ possibly zero), such that
+$X_{i}\cap\partial\Sigma_{j}=\cup_{1\leq p\leq n(i,j)}\,T_{i}\big(\{0\}\times\{\theta(i,j,p)\}\times(-1,1)\big)$ (35)
+holds, with normal vectors ${\bm{n}}_{\Sigma_{i}}$ and indices at the intersection points $P_{i,j,p}=T_{i}(0,\theta(i,j,p),0)$,
+${\rm ind}(i,j,p):={\rm sign}\Big(({\bm{\tau}}_{\Gamma_{i}}\cdot{\bm{n}}_{\Sigma_{j}})(P_{i,j,p})\Big)\in\{-1,+1\},$ (36)
+that satisfy the duality relation
+$I(\Gamma_{i},\Sigma_{j}):=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (37)
+In our analysis below we will consider this general case since it is handled by the same arguments.
+Remark 3.5.
+The fact that the curves $\Gamma_{i}$ represent a basis of the first homology group $\mathcal{H}_{1}(\overline{\Omega})$ follows from Assumption 3.3 and properties of the intersection pairing (29) recalled at the beginning of Section 3.1.
+Indeed, these curves being 1-cycles they represent 1-chain homology classes, and similarly the surfaces $\Sigma_{i}$ being closed modulo the boundary they represent relative homology classes in $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$.
+Their intersection duality assumed in (34) (or more generally in (35)–(37)) then yields $Q([\Gamma_{i}],[\Sigma_{j}])=\delta_{i,j}$, which shows the linear independence of the classes $\Gamma_{i}]$.
+Their generating property follows from the definition of $\beta_{1}$.
 3.2 The surface flux functionals
-Our assumptions on the surfaces $\Sigma_{i}$ guarantee that they are surface with corners [24], and more generally standard surfaces in the sense of [36, Sec.
-III-18], so that the Stokes theorem applies.
-Using the correspondence between vector fields and 1-forms, this yields
-$\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{u}}$ (32)
-for any vector field ${\bm{u}}\in C^{1}(\overline{\Omega})^{3}$.
-We observe that (32) also holds for ${\bm{u}}\in H(\operatorname{{\bf curl}};\Omega)$ which is piecewise $C^{1}$ in the sense specified above.
-Indeed, for every interface $f=\overline{M_{a}}\cap\overline{M_{b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, such a field satisfies
-${\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{a}}={\bm{\tau}}_{f}\cdot{\bm{u}}|_{M_{b}}\quad\text{ on }f.$ (33)
-Decomposing $\Sigma_{i}$ into subsurfaces $\Sigma_{i,a}=\Sigma_{i}\cap M_{a}$ with piecewise $C^{1}$ boundaries, one then applies the Stokes formula on each of these subsurfaces which yields a collection of subsurface integrals which sum to the surface integral in (32), and two types of line integrals: line integrals corresponding to interior curves (shared by two subsurfaces) $\gamma_{i,a,b}=\partial\Sigma_{i,a}\cap\partial\Sigma_{i,b}$ and line integrals over boundary curves $\gamma_{i,a}=\partial\Sigma_{i,a}\cap\partial\Sigma$.
-Since the former ones appear twice with opposite orientations, their sum vanishes thanks to (33), and the latter ones sum to the boundary integral in (32).
-Using the same arguments as in the proof of [1, Lemma 3.10], which rely on the properties of $\Sigma_{i}$ listed in Assumption 3.1, we can extend the mapping $\gamma_{n}:C^{0}_{c}(\Omega)\to L^{\infty}(\Sigma_{i}),{\bm{w}}\mapsto{\bm{n}}\cdot{\bm{w}}|_{\Sigma_{i}}$ to a bounded trace operator $\gamma_{n}:H_{0}(\operatorname{div};\Omega)\to(H^{1/2}(\Sigma_{i}))^{\prime}.$ For each $1\leq i\leq\beta_{1}$, this allows us to consider a flux functional on $\Sigma_{i}$ defined as
-$\Phi_{i}:H_{0}(\operatorname{div};\Omega)\to\mathbb{R},\quad{\bm{w}}\mapsto\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}.$ (34)
+The main purpose of the reciprocal surfaces $\Sigma_{i}$ is to establish the linear independence of our tangent harmonic vector fields.
+To this end we first review a version of the usual Stokes formula for piecewise smooth curl-conforming fields.
 Proposition 3.6.
-The linear form $\Phi_{i}$ vanishes on $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$, i.e., we have
-$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}})=0,\quad{\bm{v}}\in H_{0}(\operatorname{{\bf curl}};\Omega).$ (35)
+Let $\Sigma$ be an oriented surface compatible with a mesh $\mathcal{M}$ in the sense of Definition 3.2. Then the Stokes formula
+$\int_{\Sigma}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma}{\bm{\tau}}\cdot{\bm{u}}$ (38)
+holds for any ${\bm{u}}\in C^{\infty}(\mathcal{M})\cap H(\operatorname{{\bf curl}};\Omega)$.
 Proof.
-For ${\bm{v}}\in C^{\infty}_{c}(\Omega)$, Stokes’ formula (32) yields
-$\Phi_{i}(\operatorname{{\bf curl}}{\bm{v}}_{n})=\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{v}}_{n}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{v}}_{n}=0$ (36)
-where the last equality follows from $\partial\Sigma_{i}\subset\partial\Omega$.
-We conclude with a density argument: given a sequence ${\bm{v}}_{n}\in C^{\infty}_{c}(\Omega)$ converging to an arbitrary $\bm{v}$} in $H_{0}(\operatorname{{\bf curl}};\Omega)$, we have $\operatorname{{\bf curl}}{\bm{v}}_{n}\to\operatorname{{\bf curl}}{\bm{v}}$ in $H_{0}(\operatorname{div};\Omega)$ and the result follows from the continuity of the operator (34).
+Since any surface $\Sigma_{a}$, $a=1,\dots n$, involved in the decomposition of Definition 3.2 is a smooth surface with corners and is included in some domain $\overline{M}_{b(a)}$ where $\bm{u}$} is smooth, Stokes’ theorem applies.
+Using the correspondence between vector fields and differential forms (see for instance [30, Ch.
+16]), this yields
+$\int_{\Sigma_{a}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{a}}{\bm{\tau}}\cdot{\bm{u}}|_{M_{b(a)}},\qquad 1\leq a\leq n.$ (39)
+Summing over $a$ then yields a collection of surface integrals which sum to the left-hand side integral in (38), and a collection of line integrals on the right-hand side: some corresponding to interior curves $\gamma_{a,a^{\prime}}=\partial\Sigma_{a}\cap\partial\Sigma_{a^{\prime}}$ with $a\neq a^{\prime}$, and others corresponding to boundary curves $\gamma_{a}=\partial\Sigma_{a}\cap\partial\Sigma$.
+On interior curves which appear twice with opposite orientations, we have
+$\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a}}\cdot{\bm{u}}|_{M_{b(a)}}+\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a^{\prime}}}\cdot{\bm{u}}|_{M_{b(a^{\prime})}}=\int_{\gamma_{a,a^{\prime}}}{\bm{\tau}}_{\partial\Sigma_{a}}\cdot({\bm{u}}|_{M_{b(a)}}-{\bm{u}}|_{M_{b(a^{\prime})}})=0$
+where the last equality follows from the continuity of the tangential traces of piecewise smooth, globally $H(\operatorname{{\bf curl}};\Omega)$ fields, see e.g.
+[8, Prop.
+2.2.32].
+As a result, only the line integrals over boundary curves remain in the sum, yielding the right-hand side integral in (38).
+∎
+Let us further study the flux functionals defined on the surfaces $\Sigma_{i}$.
+Proposition 3.7.
+For each $1\leq i\leq\beta_{1}$, the flux functional
+$\Phi_{i}:C^{\infty}_{c}(\Omega)\to\mathbb{R},\qquad{\bm{w}}\mapsto\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}.$ (40)
+extends to a continuous linear form on $H_{0}(\operatorname{div};\Omega)$ which vanishes on $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)$: we have
+$\Phi_{i}(\operatorname{{\bf curl}}{\bm{u}})=0,\qquad{\bm{u}}\in H_{0}(\operatorname{{\bf curl}};\Omega).$ (41)
+Proof.
+The argument is similar to that used in the proof of [3, Lemma 3.10], although simpler since we consider fluxes defined on single surfaces.
+Using the fact that $\Sigma_{i}$ is Lipschitz and oriented, we can find a scalar function ${\varphi}_{i}\in H^{1}(\omega_{i})$ with $\omega_{i}=\Omega\setminus\Sigma_{i}$, whose trace vanishes on the positive side of $\Sigma_{i}$ (as pointed by its normal vector $\bm{n}$}) and is 1 on its negative side.
+For ${\bm{w}}\in C^{\infty}_{c}(\Omega)$, Green’s formula yields then
+$\Phi_{i}({\bm{w}})=\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}=\int_{\omega_{i}}{\bm{w}}\cdot\operatorname{{\bf grad}}\phi_{i}+\int_{\omega_{i}}(\operatorname{div}{\bm{w}})\phi_{i}$ (42)
+and in particular one has
+$\Phi_{i}({\bm{w}})\leq\lVert{\bm{w}}\rVert_{L^{2}(\omega_{i})}\lVert\operatorname{{\bf grad}}\phi_{i}\rVert_{L^{2}(\omega_{i})}+\lVert\operatorname{div}{\bm{w}}\rVert_{L^{2}(\omega_{i})}\lVert\phi_{i}\rVert_{L^{2}(\omega_{i})}\leq\lVert{\bm{w}}\rVert_{H(\operatorname{div};\Omega)}\lVert\phi_{i}\rVert_{H^{1}(\omega_{i})}$ (43)
+so that $\Phi_{i}$ admits indeed a continous extension in $H_{0}(\operatorname{div};\Omega)$. To verify (41) we then observe that for ${\bm{u}}\in C^{\infty}_{c}(\Omega)$, Stokes’ formula (38) gives
+$\Phi_{i}(\operatorname{{\bf curl}}{\bm{u}})=\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{u}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{u}}=0$ (44)
+where the last equality follows from the inclusion $\partial\Sigma_{i}\subset\partial\Omega$.
+We conclude by using the density of $C^{\infty}_{c}(\Omega)$ in $H_{0}(\operatorname{{\bf curl}};\Omega)$, the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$, and the continuity of $\Phi_{i}$ in the latter space.
 ∎
@@
 ${\bm{A}}_{i}:={\bm{A}}_{i}^{b}+{\bm{A}}_{i}^{0},\qquad 1\leq i\leq\beta_{1},$ (45)
-where the first terms ${\bm{A}}_{i}^{b}$ lift some suitable boundary condition into the domain, and the second terms ${\bm{A}}_{i}^{0}$ are correction terms with homogeneous boundary conditions.
-Specifically, the ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ are piecewise smooth fields (in the sense of Section 3.1) which lift the following boundary conditions:
-${\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0~~\text{ on }\partial\Omega\qquad\text{ and }\qquad\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}.$ (38)
-Notice that the first condition may be reformulated using a surface operator (see, e.g., [6, Cor.
+where the first terms lift some suitable boundary condition into the domain, and the second are correction terms with homogeneous boundary conditions.
+Specifically, the ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ are piecewise smooth fields which lift the following boundary conditions:
+${\bm{n}}_{\partial\Omega}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0~~\text{ on }\partial\Omega\qquad\text{ and }\qquad\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}$ (46)
+where ${\bm{n}}_{\partial\Omega}$ is the unit normal vector on the Lipschitz surface $\partial\Omega$, pointing outside of $\Omega$.
+Notice that the first condition may be reformulated using a surface operator (see, e.g., [8, Cor.
 3.1.16]) so that it only involves boundary values of ${\bm{A}}^{b}_{i}$ indeed.
@@
 Proposition 3.8.
-If (38) and (39) hold, then the potentials ${\bm{A}}_{i}$ satisfy $\operatorname{{\bf curl}}{\bm{A}}_{i}\in\mathfrak{H}^{1}$ and
+If (46) and (47) hold, then the potentials (45) satisfy $\operatorname{{\bf curl}}{\bm{A}}_{i}\in\mathfrak{H}^{1}$ and
 $\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i})=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (48)
@@
 The boundary condition ${\bm{w}}_{i}\in H_{0}(\operatorname{div};\Omega)$ is also easily verified, indeed it is satisfied by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}$ thanks to (46), and by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{0}$ thanks to the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$.
-This shows that the fields ${\bm{w}}_{i}$ are indeed tangent-harmonic.
+This shows that the fields ${\bm{w}}_{i}$ are indeed tangent and harmonic.
 To show the relations (48) we first consider the lifted boundary part.
-Since it is piecewise smooth, Stokes’ formula (32) yields
+Since it is piecewise smooth over the mesh $\mathcal{X}_{i}$, Stokes’ formula (38) yields
 $\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i}^{b})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
 where the last equality is (46).
-The desired relation (40) then follows from the fact that the flux functionals $\Phi_{j}$ vanish for the ${\bm{A}}^{0}_{i}$ parts, thanks to (35).
-The linear independence of the fields ${\bm{w}}_{i}$’s follows from (40), and the fact that they span $\mathfrak{H}^{1}$ follows from (10).
+The desired relation (48) then follows from the fact that the flux functionals $\Phi_{j}$ vanish for the ${\bm{A}}^{0}_{i}$ parts thanks to (41).
+The linear independence of the fields ${\bm{w}}_{i}$ readily follows, and the fact that they span $\mathfrak{H}^{1}$ is a consequence of de Rham’s Theorem 2.1.
 ∎
@@
 $\hat{{\bm{n}}}\cdot\operatorname{{\bf curl}}\widehat{{\bm{A}}}^{b}=0\quad\text{ on }r=0,$ (49)
-where $\hat{{\bm{n}}}=-\hat{{\bm{e}}}_{r}$ is the unit normal vector pointing outside of $\mathcal{O}$ on its $\{r=0\}$ boundary, and
-$\int_{-1}^{1}\hat{\bm{\tau}}\cdot\widehat{{\bm{A}}}^{b}(0,0,z)\ \mathrm{d}z=1$ (42)
-where $\hat{\bm{\tau}}=\hat{{\bm{e}}}_{z}$ is the unit vector tangent to the $\{r=0,\theta=0\}$ curve.
-Since $\widehat{{\bm{A}}}^{b}$ is compactly supported, it vanishes on every boundary of $\mathcal{O}$ with the exception of the surface $\{r=0\}$ which is contained in $\mathcal{O}$: it is on this boundary that the above boundary conditions are imposed.
+where $\hat{{\bm{n}}}=-\hat{{\bm{e}}}_{r}$ is the unit normal vector pointing outside of $\mathcal{O}$ on its $\{r=0\}$ boundary (see Figure 2), and
+$\int_{-1}^{1}\hat{{\bm{e}}}_{z}\cdot\widehat{{\bm{A}}}^{b}(0,\theta,z)\ \mathrm{d}z=1,\qquad\forall\theta\in\mathbb{T}.$ (50)
+Note that since $\widehat{{\bm{A}}}^{b}$ is compactly supported in $\mathcal{O}$, it vanishes on every face with the exception of the face $\{r=0\}$ which is contained in $\mathcal{O}$: it is on this face that the above boundary conditions are imposed.
 Specifically, we may set
@@
 where $\mathcal{F}^{1}_{i}(\hat{\bm{u}}):=((DT_{i})^{-T}\hat{\bm{u}})\circ T_{i}^{-1}$ is the covariant Piola transform (with $DT_{i}$ the Jacobian matrix of $T_{i}$) corresponding to the pushforward operator for 1-form proxy fields [28].
-Let us denote by $\mathcal{O}_{i,a}$, $a=1,\dots,m_{i}$ the subdomains of $\mathcal{O}$ on which $T_{i}$ is smooth.
-On the corresponding domains $X_{i,a}=T_{i}(\mathcal{O}_{i,a})$ the field ${\bm{A}}_{i}^{b}$ is smooth as well.
 Proposition 3.9.
-The field ${\bm{A}}_{i}^{b}$ just defined satisfies the properties stated above, namely it is $H(\operatorname{{\bf curl}};\Omega)$ and the boundary conditions (38) hold.
+The field ${\bm{A}}_{i}^{b}$ just defined satisfies the properties stated above: it is piecewise $C^{\infty}$ over the mesh $\mathcal{X}_{i}$ in (33), it belongs to $H(\operatorname{{\bf curl}};\Omega)$ and the boundary conditions (46) hold.
 Proof.
-To verify that ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ we need to show that (i) the tangential trace of the pushforward $\mathcal{F}_{i}^{1}(\widehat{{\bm{A}}}^{b})({\bm{x}})$ on $\partial X_{i}\setminus\partial\Omega$ vanishes, and (ii) on every interface $f=\overline{X_{i,a}}\cap\overline{X_{i,b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, we have
+By construction ${\bm{A}}_{i}^{b}$ vanishes on $\overline{X}_{i,0}$, and its smoothness on $\overline{X}_{i,a}$, $a=1,\dots m_{i}$, follows from the smoothness of the logical potential (51) and that of the mapping $T_{i}$ and its inverse on the subdomains $\overline{\mathcal{O}}_{i,a}$ and $\overline{X}_{i,a}=T_{i}(\overline{\mathcal{O}}_{i,a})$.
+To verify that ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$, according to [8, Prop.
+2.2.32] we need to show that (i) the tangential trace of the potential (52) vanishes on $\partial X_{i}\setminus\partial\Omega$, and (ii) on every interface $f=\overline{X}_{i,a}\cap\overline{X}_{i,b}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, we have
 ${\bm{\tau}}_{f}\cdot{\bm{A}}_{i}^{b}|_{X_{i,a}}={\bm{\tau}}_{f}\cdot{\bm{A}}_{i}^{b}|_{X_{i,b}}\quad\text{ on }f.$ (53)
-The former property is clear since the logical field (43) vanishes on a neighborhood of $\partial\mathcal{O}\setminus\{r=0\}$, and the latter property follows from from the fact that a tangent vector on $f$ is of the form
+The former property is clear since the field (51) vanishes on a neighborhood of $\partial\mathcal{O}\setminus\{r=0\}$, and the latter property follows from from the fact that a tangent vector on $f$ is of the form
 […]
-where $\hat{\bm{\tau}}$ is a vector tangent to the logical surface $\hat{f}=T_{i}^{-1}(f)$, and where the Jacobian matrix of $T_{i}$ can be evaluated from both subdomains $\mathcal{O}_{i,a}$ or $\mathcal{O}_{i,b}$ since the directional derivative is taken along their common interface.
+where $\hat{\bm{\tau}}$ is a vector tangent to the surface $\hat{f}=T_{i}^{-1}(f)$, and where the Jacobian matrix of $T_{i}$ can be evaluated from both subdomains $\mathcal{O}_{i,a}$ or $\mathcal{O}_{i,b}$ since the directional derivative is taken along their common interface.
 The continuity of the tangential traces on $f$ then follows from the definition of the pushforward.
-Finally to verify the boundary conditions we use the commutation relation
+We finally turn to the boundary conditions (46). Using the commutation relation
 $\mathcal{F}^{2}_{i}\operatorname{{\bf curl}}\widehat{{\bm{A}}^{b}}=\operatorname{{\bf curl}}\mathcal{F}^{1}_{i}\widehat{{\bm{A}}^{b}}$
 satisfied by the 2-form pushforward operator $\mathcal{F}^{2}_{i}$, see e.g.
-[22].
-The relations (38) then follow from (41), (42), and the fact that line and surface integrals of smooth vector fields are respectively preserved by 1-form and 2-form pushforward operators [24, Ch.
-16.6].
-∎
+[28], and the fact that surface integrals of smooth vector fields are preserved by 2-form pushforwards [30, Ch.
+16.6], we verify that the first boundary condition ${\bm{n}}_{\partial\Omega}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0$ is satisfied as a consequence of its parametric counterpart (49).
+To verify the second one, we remind that the intersection of $\partial\Sigma_{j}$ with $X_{i}$ decomposes into $n(i,j)$ disjoint curves according to (35)–(37) (or (34) in the simpler case where $n(i,j)=\delta_{i,j}$) that is,
+$X_{i}\cap\partial\Sigma_{j}=\cup_{1\leq p\leq n(i,j)}\tilde{\Gamma}(i,j,p)\quad\text{ with }~\tilde{\Gamma}(i,j,p)=T_{i}\big(\{0\}\times\{\theta(i,j,p)\}\times(-1,1)\big)$
+each curve being oriented by a unit vector induced by the global orientation of $\Sigma_{j}$ as defined by its normal vector field ${\bm{n}}_{\Sigma_{j}}$.
+Using the fact that line integrals are preserved by 1-form pushforwards, we then write
+$\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}\int_{\tilde{\Gamma}(i,j,p)}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}\int_{-1}^{1}\hat{\bm{\tau}}(i,j,p)\cdot\widehat{{\bm{A}}_{i}^{b}}(0,\theta(i,j,p),z)\ \mathrm{d}z$
+where $\hat{\bm{\tau}}(i,j,p)$ is the tangent vector of the straight curve $T_{i}^{-1}(\tilde{\Gamma}(i,j,p))$ in the direction induced by that of $\tilde{\Gamma}(i,j,p)$, that is,
+$\hat{\bm{\tau}}(i,j,p)={\rm ind}(i,j,p)\hat{{\bm{e}}}_{z}$
+according to (36). Using (50) and (37) then yields the desired result:
+$\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)\int_{-1}^{1}\hat{{\bm{e}}}_{z}\cdot\widehat{{\bm{A}}_{i}^{b}}(0,\theta(i,j,p),z)\ \mathrm{d}z=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)=\delta_{i,j}.$
 3.5 Correction potentials
@@
 Proposition 3.10.
-The system (46) defines a unique field ${\bm{A}}_{i}^{0}$ which satisfies (39).
+System (54) defines a unique field ${\bm{A}}_{i}^{0}$ which satisfies (47).
 Proof.
-The well-posedness follows by applying [5, Theorem 3.2], to the adjoint (homogeneous) de Rham complex (2), which is possible thanks to the properties stated in [5, Sec.
+The well-posedness follows by applying [7, Theorem 3.2], to the adjoint (homogeneous) de Rham complex (2), which is possible thanks to the properties stated in [7, Sec.
 4.2].
@@
 ∎
-Combining Propositions 3.4, 3.5 and 3.6 yields the following result.
+3.6 Main result and invariance principles
+By combining Propositions 3.8, 3.9 and 3.10, which hold under Assumption 3.3, we have proven our main result.
 Theorem 3.1.
-The fields ${\bm{A}}_{i}={\bm{A}}^{b}_{i}+{\bm{A}}^{0}_{i}$ obtained with the above construction are vector potentials for the tangent harmonic fields.
-In particular, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}$ that is dual to the flux functionals $\Phi_{i}$.
+The fields ${\bm{A}}_{i}={\bm{A}}^{b}_{i}+{\bm{A}}^{0}_{i}$ constructed over the tunnel curves $\Gamma_{i}$, $i=1,\dots\beta_{1}$ as described above, are free vector potentials for the tangent harmonic fields.
+More precisely, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$ form a basis for the space $\mathfrak{H}^{1}$ which is dual to the flux functionals $\Phi_{i}$ associated with the reciprocal surfaces, in the sense that
+$\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}=\delta_{i,j}$ (55)
+holds for all $1\leq i,j\leq\beta_{1}$.
+Our construction satisfies two (dual) invariance principles.
+The first one is with respect to the tunnel curves: if $\Gamma_{i}$ and $\Gamma^{\prime}_{i}$, $i=1,\dots\beta_{1}$, are two collections of piecewise smooth closed boundary curves satisfying the properties of Assumption 3.3 with the same reciprocal surfaces $\Sigma_{i}$, then their associated harmonic potentials are a priori different but their harmonic fields coincide
+$\operatorname{{\bf curl}}{\bm{A}}_{i}=\operatorname{{\bf curl}}{\bm{A}}^{\prime}_{i},\quad i=1,\dots\beta_{1},$ (56)
+since it follows from Theorem 3.1 that the fluxes $\Phi_{i}$ are unisolvent on the harmonic space $\mathfrak{H}^{1}$.
+This invariance principle actually holds with respect to the 1-chain homology classes: indeed if $\Gamma_{i}$ and $\Gamma^{\prime}_{i}$ are two collections of piecewise smooth closed boundary curves that represent the same basis $\Gamma_{i}]$ of the first homology group $\mathcal{H}_{1}(\overline{\Omega})$, then we may consider the (unique) basis $\Sigma_{i}]$ of the relative homology group $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$ which is dual in the sense of the intersection pairing (29), and any collection of piecewise smooth surfaces $\Sigma_{i}$ which represent this dual basis while intersecting nicely all the tunnel curves.
+Because the intersection pairing coincides with the algebraic intersection number of any pair of representative cycles, we find that the surfaces $\Sigma_{i}$ are reciprocal to both curve collections, so that the associated harmonic fields coincide again.
+A dual invariance principle holds with respect to the reciprocal surfaces: indeed if $\Sigma_{i}$ and $\Sigma^{\prime}_{i}$, $i=1,\dots\beta_{1}$, are two collections of piecewise smooth surfaces wich are reciprocal to the same tunnel curves $\Gamma_{i}$ in the sense of Assumption 3.3, then the harmonic fields ${\bm{w}}_{i}=\operatorname{{\bf curl}}{\bm{A}}_{i}$ associated with these curves have the same fluxes across both surface collections:
+$\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}_{i}=\int_{\Sigma^{\prime}_{i}}{\bm{n}}\cdot{\bm{w}}_{i},\quad i=1,\dots\beta_{1},$ (57)
+as an immediate consequence of Theorem 3.1.
+Again, it follows from the properties of the intersection pairing (29) that this dual invariance principle actually holds with respect to the homology classes of 2-chains modulo the boundary.
 4 Structure-preserving discrete harmonic potentials
 We end this article by describing how the above construction can be implemented in a discrete setting.
-The approach proposed here applies to structure-preserving finite element spaces that share a certain number of properties listed in Sections 4.1 and 4.3 below.
-Assuming that the domain $\Omega$ can be triangulated with a simplicial mesh, these properties hold for instance in the case of Whitney [36, 9] and Nédélec [29] finite element spaces, and more generally for high order FEEC spaces [4], but they are not restricted to these cases.
-4.1 Discretization of the homogeneous (dual) and extended dual complexes
+The approach proposed here applies to structure-preserving finite element spaces that share a certain number of properties listed in Sections 4.1 and 4.2 below.
+Assuming that the domain $\Omega$ can be triangulated with a simplicial mesh, these properties hold for instance in the case of Whitney [44, 12] and Nédélec [37] finite element spaces, and more generally high order FEEC spaces [6], but they are not restricted to these cases.
+4.1 Discretization of the homogeneous and extended dual complexes
 Since harmonic fields can be described by potentials, using differential operators from the dual complex extended to spaces without homogeneous boundary conditions, a natural choice is to consider a strong discretization of the dual de Rham complex (2), seen as a subcomplex of the one without boundary conditions (i.e., the primal one).
@@
 For simplicity we may disregard these sign changes and consider the plain differential operators below.
-However, we should not confuse the primal and the extended dual complexes.
-We then consider a discretization of the extended dual complex equipped with degrees of freedom that are commuting and compatible with the Dirichlet boundary conditions.
-By this we mean essentially two properties.
-First, that we are given a discrete subcomplex of (48),
-$0\xleftarrow{~~}\tilde{V}^{3}_{h}\xleftarrow{~\operatorname{div}~}\tilde{V}^{2}_{h}\xleftarrow{~\operatorname{{\bf curl}}~}\tilde{V}^{1}_{h}\xleftarrow{~\operatorname{{\bf grad}}~}\tilde{V}^{0}_{h}\xleftarrow{~~}0$ (49)
-with spaces of piecewise continuous functions $\tilde{V}^{\ell}_{h}\subset C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ (associated with some mesh $\mathcal{M}_{h}$ of possibly curved or polygonal Lipschitz cells, as described in Section 3.1) equipped with linear degrees of freedom
-$\tilde{\sigma}^{\ell}_{i}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\mathbb{R},\qquad 1\leq i\leq N_{\ell}:=\dim(\tilde{V}^{\ell}_{h}),$ (50)
-satisfying the following commuting relation: for all $v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ with $\tilde{d}^{\ell}v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell+1}$, it holds
-$\tilde{\sigma}^{\ell+1}_{i}(\tilde{d}^{\ell}v)=\sum_{j=1}^{N_{\ell}}\tilde{\mathbb{D}}^{\ell}_{ij}\tilde{\sigma}^{\ell}_{j}(v),\qquad 1\leq i\leq N_{\ell+1},$ (51)
-with a fixed $N_{\ell+1}\times N_{\ell}$ matrix $\mathbb{D}^{\ell}$ that represents the differential operator at the discrete level.
-We note that a relation of the form (51) actually implies that the corresponding projection operators $\tilde{\Pi}^{\ell}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\tilde{V}^{\ell}_{h}$, defined by
-$\tilde{\sigma}^{\ell}_{i}(\tilde{\Pi}^{\ell}v)=\tilde{\sigma}^{\ell}_{i}(v),\qquad 1\leq i\leq N_{\ell},$ (52)
-commute with the differential operators. The second property is that the degrees of freedom lead to projection operators that preserve the homogenous boundary conditions, in the sense that we have
-$\tilde{\Pi}^{\ell}(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}_{0})=(\tilde{V}^{\ell}_{h}\cap\tilde{V}^{\ell}_{0})=:\tilde{V}^{\ell}_{h,0}.$ (53)
-In particular, the same projection operators yield two commuting diagrams: one for the dual complex (with homogeneous boundary conditions)
-[…]
-and another one for the extended dual complex,
-[…]
-4.2 Flux functionals on discrete surfaces
-To study the linear independence of our discrete vector potentials, we assume in addition that there exist surfaces $\Sigma_{h,i}$ made of smooth surface with corners, with the following properties: (i) the line integrals over the boundaries $\partial\Sigma_{h,i}$ can be expressed as linear combinations of degrees of freedom in $\tilde{V}^{1}_{h}$, (ii) these boundaries are included in the domain boundary ($\partial\Sigma_{h,i}\subset\partial\Omega$) and (iii) they coincide with those of the smooth reciprocal surfaces $\Sigma_{i}$ on the parametrized domains $X_{j}$, in the sense that
-$X_{j}\cap\partial\Sigma_{h,i}=X_{j}\cap\partial\Sigma_{i},\qquad 1\leq i,j\leq\beta_{1}.$ (56)
-We remind that these intersections are empty for $i\neq j$, see (31).
-In the case of Whitney, Nédélec or more general FEEC finite elements spaces, the usual degrees of freedom involve line integrals along edges of the mesh, so that the first property holds for surfaces $\Sigma_{h,i}$ made of faces of the mesh $\mathcal{M}_{h}$.
-The second property poses no particular difficulty, and the last one can then be fulfilled by choosing a parametrization of $\Gamma_{i}$ such that the smooth curves $X_{i}\cap\partial\Sigma_{i}$ are included in a single edge of the boundary $\partial\Sigma_{h,i}$.
-In turn, this implies that the finite element projection $\tilde{\Pi}^{1}$ interpolates the corresponding line integrals: indeed one has
-$\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}(\tilde{\Pi}^{1}{\bm{a}})=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}({\bm{a}})=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$
-for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$, where the coefficients $c_{i,j}$ correspond to the decomposition of the line integrals in terms of degrees of freedom.
-Moreover, applying Stokes’ theorem on each smooth face $f\in F(\Sigma_{h,i})$, we compute
-$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\sum_{f\in F(\Sigma_{h,i})}\int_{f}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{a}}_{h},\qquad{\bm{a}}_{h}\in\tilde{V}^{1}_{h},$
-where the last equality uses the fact that integrals on interior curves, i.e., curves of the type $\partial f\cap\partial f^{\prime}$ with $f\neq f^{\prime}$, cancel because of their opposite orientation and the tangential continuity of the fields in $H(\operatorname{{\bf curl}};\Omega)$.
-As a consequence, we obtain
-$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=0$ (57)
-for all ${\bm{a}}_{h}\in\tilde{V}^{1}_{0,h}$, and
-$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$ (58)
-for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$.
+We distinguish, however, the primal and the extended dual complexes which play different roles.
+We then consider a discrete subcomplex of (59),
+$0\xleftarrow{~~}\tilde{V}^{3}_{h}\xleftarrow{~\operatorname{div}~}\tilde{V}^{2}_{h}\xleftarrow{~\operatorname{{\bf curl}}~}\tilde{V}^{1}_{h}\xleftarrow{~\operatorname{{\bf grad}}~}\tilde{V}^{0}_{h}\xleftarrow{~~}0,$ (60)
+made of finite-dimensional spaces of piecewise $C^{\infty}$ functions over some mesh $\mathcal{M}_{h}$ of $\Omega$, made of cells with Lipschitz and piecewise smooth boundaries as in Definition 3.1.
+Thus, we consider discrete spaces
+$\tilde{V}^{\ell}_{h}\subset C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell},\qquad 0\leq\ell\leq 3,$ (61)
+with the following properties:
+  • (P1)
+    $\mathcal{M}_{h}$ is finer than every mesh $\mathcal{X}_{i}$ from (33), in the sense that for all $1\leq i\leq\beta_{i}$ it holds
+    $C^{\infty}(\mathcal{X}_{i})\subseteq C^{\infty}(\mathcal{M}_{h})$ (62)
+    moreover $\mathcal{M}_{h}$ is compatible with every surface $\Sigma_{i}$ in the sense of Definition 3.2,
+  • (P2)
+    each space $\tilde{V}^{\ell}_{h}$ is equipped with degrees of freedom
+    $\tilde{\sigma}^{\ell}_{i}:(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\mathbb{R},\qquad 1\leq i\leq\tilde{N}_{\ell}:=\dim(\tilde{V}^{\ell}_{h}),$ (63)
+    satisfying relations of the form
+    $\tilde{\sigma}^{\ell+1}_{i}(\tilde{d}^{\ell}v)=\sum_{j=1}^{\tilde{N}_{\ell}}\tilde{\mathbb{D}}^{\ell}_{ij}\tilde{\sigma}^{\ell}_{j}(v),\qquad 1\leq i\leq\tilde{N}_{\ell+1},$ (64)
+    for all $v\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ with $\tilde{d}^{\ell}v\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell+1}$, and some $\tilde{N}_{\ell+1}\times\tilde{N}_{\ell}$ matrices $\tilde{\mathbb{D}}^{\ell}$.
+The matrices $\tilde{\mathbb{D}}^{\ell}$ represent the differential operators $\tilde{d}^{\ell}$ at the discrete level, and we note that the relations (64) imply that the corresponding projection operators $\tilde{\Pi}^{\ell}:(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\tilde{V}^{\ell}_{h}$, defined by
+$\tilde{\sigma}^{\ell}_{i}(\tilde{\Pi}^{\ell}v)=\tilde{\sigma}^{\ell}_{i}(v),\qquad 1\leq i\leq\tilde{N}_{\ell}$, (65)
+commute with the differential operators. Further, we assume that
+  • (P3)
+    the projection operators preserve the homogenous boundary conditions, in the sense that
+    $\tilde{\Pi}^{\ell}(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}_{0})=\tilde{V}^{\ell}_{h}\cap\tilde{V}^{\ell}_{0}=:\tilde{V}^{\ell}_{h,0}$ (66)
+  • (P4)
+    the line integrals over the reciprocal surface boundaries $\partial\Sigma_{i}$, $1\leq i\leq\beta_{1}$, can be expressed as linear combinations of degrees of freedom in $\tilde{V}^{1}_{h}$.
+We observe that in practice Property (P1) is not very restrictive since in most cases one can subdivide the surfaces $\Sigma_{i}$ into smaller pieces to make them compatible with a fine mesh.
+In the case of Whitney, Nédélec or more general FEEC finite elements, the usual degrees of freedom satisfy property (P2) and (P3).
+Moreover the curl-conforming degrees of freedom $\tilde{\sigma}^{1}_{i}$ involve line integrals along edges of the mesh, so that property (P4) holds as long as one can build the reciprocal surfaces $\Sigma_{i}$ by assembling some faces of the mesh $\mathcal{M}_{h}$ (which readily makes them compatible).
+It follows from properties (P2) and (P3) that the same projection operators yield two commuting diagrams: one for the dual complex (58) (with homogeneous boundary conditions)
+$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{3}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{2}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{0}_{0}~\phantom{\big|}$$\tilde{V}^{3}_{h,0}~\phantom{\big|}$$\tilde{V}^{2}_{h,0}~\phantom{\big|}$$\tilde{V}^{1}_{h,0}~\phantom{\big|}$$\tilde{V}^{0}_{h,0}~\phantom{\big|}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\tilde{\Pi}^{3}$$\tilde{\Pi}^{2}$$\tilde{\Pi}^{1}$$\tilde{\Pi}^{0}$ (67)
+and another one for the extended dual complex (59),
+$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{3}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{2}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{0}~\phantom{\big|}$$\tilde{V}^{3}_{h}~\phantom{\big|}$$\tilde{V}^{2}_{h}~\phantom{\big|}$$\tilde{V}^{1}_{h}~\phantom{\big|}$$\tilde{V}^{0}_{h}~.\phantom{\big|}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\tilde{\Pi}^{3}$$\tilde{\Pi}^{2}$$\tilde{\Pi}^{1}$$\tilde{\Pi}^{0}$ (68)
+Our assumptions also imply that the projection $\tilde{\Pi}^{1}$ interpolates the line integrals along the boundaries $\partial\Sigma_{i}$, $1\leq i\leq\beta_{1}$: one has indeed
+$\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\sum_{j=1}^{\tilde{N}_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}(\tilde{\Pi}^{1}{\bm{a}})=\sum_{j=1}^{\tilde{N}_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}({\bm{a}})=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{a}},\qquad{\bm{a}}\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1},$
+where the coefficients $c_{i,j}$ correspond to the decomposition of the line integrals in terms of degrees of freedom. In particular, Stokes’ formula (38) yields
+$\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{a}},\qquad{\bm{a}}\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}.$ (69)
 4.2 Discrete harmonic spaces
-A discretization of the harmonic spaces (5) is classically provided by the harmonic spaces associated with the discrete complex [4].
-Since the primal and dual complexes have the same harmonic spaces at the continuous level, one can use the dual discrete complex (54): this leads to the discrete vector-valued spaces
+A discretization of the continuous harmonic spaces (5) is classically provided by the harmonic spaces of the discrete complex [6].
+Since the primal and dual complexes have the same harmonic spaces at the continuous level, one may use the dual discrete complex (67): this leads to the discrete vector-valued spaces
 $\tilde{\mathfrak{H}}^{2}_{h,0}:=\tilde{\mathfrak{Z}}^{2}_{h,0}\cap(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp},\qquad\tilde{\mathfrak{H}}^{1}_{h,0}:=\tilde{\mathfrak{Z}}^{1}_{h,0}\cap(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0})^{\perp},$ (70)
-where the discrete kernel spaces are defined as usual,
+where the discrete kernels are defined as usual,
 $\tilde{\mathfrak{Z}}^{2}_{h,0}:=\{{\bm{w}}\in\tilde{V}^{2}_{h,0}:\operatorname{div}{\bm{w}}=0\},\qquad\tilde{\mathfrak{Z}}^{1}_{h,0}:=\{{\bm{v}}\in\tilde{V}^{1}_{h,0}:\operatorname{{\bf curl}}{\bm{v}}=0\}.$ (71)
 Given the relabeling $\tilde{V}^{\ell}_{0}=V^{*}_{3-\ell}$, we note that the spaces (70) provide a discretization for the continuous harmonic spaces $\mathfrak{H}^{1}$ and $\mathfrak{H}^{2}$ respectively, so that we may also denote them as
-$\mathfrak{H}^{1}_{h}:=\tilde{\mathfrak{H}}^{2}_{h,0},\qquad\mathfrak{H}^{2}_{h}:=\tilde{\mathfrak{H}}^{1}_{h,0}.$ (61)
+$\mathfrak{H}^{1}_{h}:=\tilde{\mathfrak{H}}^{2}_{h,0},\qquad\mathfrak{H}^{2}_{h}:=\tilde{\mathfrak{H}}^{1}_{h,0}.$
+Remark 4.1.
+The choice made here of discretizing the tangent harmonic spaces $\mathfrak{H}^{1}_{h}$ with 2-forms is consistent with the use of (strong) vector potentials discretized as 1-forms.
+A different choice is made in [1], where the harmonic fields of $\mathfrak{H}^{1}$ are represented by discrete loop fields, which are discrete 1-forms with nonzero circulations along 1-cycles that correspond to our tunnel curves.
 We assume from now on that these discrete harmonic spaces have the same dimensions as their continuous counterparts, namely
 $\dim(\mathfrak{H}^{1}_{h})=\beta_{1},\qquad\dim(\mathfrak{H}^{2}_{h})=\beta_{2}.$ (72)
-This property holds for Whitney, Nedelec and general FEEC spaces [4, Sec. 5.5]. Finally, we can verify that discrete potentials may a priori be used for the discrete harmonic fields.
+This property holds for Whitney, Nedelec and general FEEC spaces [6, Sec. 5.5]. Finally, we can verify that discrete potentials may a priori be used for the discrete harmonic fields.
 Proposition 4.2.
@@
 We have ${\bm{v}}_{h}\in H_{0}(\operatorname{{\bf curl}}0;\Omega)$ by construction and the argument used to show (21) yields ${\bm{v}}_{h}=\operatorname{{\bf grad}}\phi$ for some $\phi\in H^{1}(\Omega)$.
-Since ${\bm{v}}_{h}$ is bounded as a piecewise continuous function, we observe that $\phi$ is also (Lipschitz) continuous, in particular we may apply the commuting projecion operators (55).
+Since ${\bm{v}}_{h}$ is piecewise smooth (and bounded), we observe that $\phi$ is also piecewise smooth, in particular we may apply the commuting projecion operators (68).
 This yields
@@
 Here we may consider an explicit potential defined on a cube $C_{R}=(-R,R)^{3}$ containing $\overline{\Omega}$,
-$\bar{\bm{A}}({\bm{x}})=\Big{(}0,\int_{-R}^{x_{1}}\bar{w}_{h,3}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime},-\int_{-R}^{x_{1}}\bar{w}_{h,2}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime}\Big{)},$
+$\bar{\bm{A}}({\bm{x}})=\Big(0,\int_{-R}^{x_{1}}\bar{w}_{h,3}(x^{\prime},x_{2},x_{3})\ \mathrm{d}x^{\prime},-\int_{-R}^{x_{1}}\bar{w}_{h,2}(x^{\prime},x_{2},x_{3})\ \mathrm{d}x^{\prime}\Big),$
 where $\bar{\bm{w}}_{h}$ is defined as the extension of ${\bm{w}}_{h}$ by 0 on $C_{R}\setminus\Omega$.
-From ${\bm{w}}_{h}\in C^{0}(\mathcal{M}_{h})$ we infer that ${\bm{A}}\in C^{0}(\mathcal{M}_{h})\cap H(\operatorname{{\bf curl}};\Omega)$, so that we may again apply the commuting projection operators (55).
+From ${\bm{w}}_{h}\in C^{\infty}(\mathcal{M}_{h})$ we infer that ${\bm{A}}\in C^{\infty}(\mathcal{M}_{h})\cap H(\operatorname{{\bf curl}};\Omega)$, so that we may again apply the commuting projection operators (68).
 This yields
@@
 Theorem 4.1.
-The fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$, $i=1,\dots,\beta_{2}$, form a basis of $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$. Moreover, they satisfy
+The fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$, $i=1,\dots\beta_{2}$, form a basis of $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$. Moreover, they satisfy
 $\langle\operatorname{{\bf grad}}\psi_{h,j},{\bm{v}}_{h,i}\rangle=\delta_{i,j},\quad 1\leq i,j\leq\beta_{2}.$ (77)
@@
 Find $(\sigma_{h,i},{\bm{A}}^{0}_{h,i},{\bm{p}}_{h,i})\in\tilde{V}^{0}_{h,0}\times\tilde{V}^{1}_{h,0}\times\tilde{\mathfrak{H}}^{1}_{h,0}$, such that
-$\left\{\begin{aligned} \langle\sigma_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}},,\rangle\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\tilde{\mathfrak{H}}^{1}_{h,0}.\end{aligned}\right$.
+$\left\{\begin{aligned} \langle\sigma_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}}\rangle,\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\tilde{\mathfrak{H}}^{1}_{h,0}.\end{aligned}\right$.
 (84)
 The well-posedness of this problem follows from the existence of uniform discrete Poincaré inequalities and bounded cochain projections, which are known to exist in the case of FEEC discretizations [6].
-Observing that the flux functionals $\Phi_{i}$ from (34) are well-defined on $\tilde{\mathfrak{H}}^{2}_{h,0}$ since the latter is a subspace of $\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)$, we then have the following result.
+We next observe that the flux functionals $\Phi_{i}$ from (40) are well-defined on $\tilde{\mathfrak{H}}^{2}_{h,0}$ since the latter is a subspace of $\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)$.
+We then have the following result.
 Theorem 4.2.
-The fields ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$.
-Moreover, their fluxes on the surfaces introduced in Section 4.2 satisfy
-$\int_{\Sigma_{h,j}}{\bm{n}}\cdot{\bm{w}}_{h,i}=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (75)
+The fields ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$, $i=1,\dots\beta_{1}$, form a basis of $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$.
+Moreover, their fluxes on the surfaces $\Sigma_{j}$ satisfy
+$\Phi_{j}({\bm{w}}_{h,i})=\int_{\Sigma_{j}}{\bm{n}}\cdot{\bm{w}}_{h,i}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}.$ (85)
 Proof.
-The proof is now a matter of simple verifications.
-By taking ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{h,i}$ in (74) we find that $\sigma_{h,i}=0$, and with ${\bm{v}}={\bm{p}}_{h,i}$ we obtain ${\bm{p}}_{h,i}=0$.
-The first and last equations from (74) yield then
-${\bm{A}}_{h,i}^{0}\in(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\overset{\perp}{\oplus}\tilde{\mathfrak{H}}^{1}_{h,0})^{\perp}=(\tilde{\mathfrak{Z}}^{1}_{h,0})^{\perp}$ (76)
-(where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (59)–(60)), while the second equation from (74) gives
-$\langle\operatorname{{\bf curl}}({\bm{A}}^{0}_{h,i}+{\bm{A}}^{b}_{h,i}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}.$ (77)
-In particular, ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$ satisfies $\operatorname{div}{\bm{w}}_{h,i}=0$ and ${\bm{w}}_{h,i}\in(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp}$.
+From (82), it is clear that ${\bm{w}}_{h,i}=\operatorname{{\bf curl}}{\bm{A}}_{h,i}\in\tilde{V}^{2}_{h}$ satisfies $\operatorname{div}{\bm{w}}_{h,i}=0$.
 As for the boundary condition ${\bm{n}}\cdot{\bm{w}}_{h,i}=0$ on $\partial\Omega$, it is verified by writing
@@
 where we have used (83) and the commutation property (68) in the second equality.
-Here both term satisfies the boundary condition: the first one thanks to (38) and (53), and the second one as the curl of a field in $H_{0}(\operatorname{{\bf curl}};\Omega)$.
-Thus we have verified that the fields ${\bm{w}}_{h,i}$ all belong to $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$ indeed.
-To evaluate the fluxes across the surfaces $\Sigma_{h,j}$, we use (57) and (58): this yields
-$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{h,i}=\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
-where in the last two steps we have used the fact that the curves $\partial\Sigma_{h,j}$ and $\partial\Sigma_{j}$ coincide on the support of ${\bm{A}}^{b}_{i}$, see (56), and the boundary conditions (38).
-This shows (75) and the linear independence of the fields ${\bm{w}}_{h,i}$.
-Their spanning properties follow from (62).
-∎
+Here both term satisfies the boundary condition: the first one thanks to (46) and (66), and the second one as the curl of a field in $\tilde{V}^{1}_{0}=H_{0}(\operatorname{{\bf curl}};\Omega)$.
+Thus, the fields ${\bm{w}}_{h,i}$ are in $\tilde{\mathfrak{Z}}^{2}_{h,0}$.
+To verify that they are in $\tilde{\mathfrak{H}}^{2}_{h,0}$ we must show that they also belong to $(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp}$, and for this we use the second equation from (84): taking first ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{h,i}$ shows that $\sigma_{h,i}=0$, then ${\bm{v}}={\bm{p}}_{h,i}$ yields ${\bm{p}}_{h,i}=0$, so that one finally obtains
+$\langle{\bm{w}}_{h,i},\operatorname{{\bf curl}}{\bm{v}}\rangle=\langle\operatorname{{\bf curl}}({\bm{A}}^{0}_{h,i}+{\bm{A}}^{b}_{h,i}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}$ (86)
+which shows that ${\bm{w}}_{h,i}\in\tilde{\mathfrak{H}}^{2}_{h,0}$ indeed.
+Finally, to evaluate the fluxes across a surface $\Sigma_{j}$ we use the fact that ${\bm{A}}^{0}_{h,i}\in H_{0}(\operatorname{{\bf curl}};\Omega)$, which yields $\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}^{0}_{h,i})=0$ according to Proposition 3.7, and (69) with ${\bm{a}}={\bm{A}}^{b}_{i}$: this gives
+$\Phi_{j}({\bm{w}}_{h,i})=\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i})=\Phi_{j}(\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{A}}^{b}_{i})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
+where the last step is again (46). This shows (85) and the linear independence of the fields ${\bm{w}}_{h,i}$. Their spanning properties follow from (72). ∎
+Remark 4.4.
+In the proof above, we further note that the first and last equations from (84) yield
+${\bm{A}}_{h,i}^{0}\in(\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\overset{\perp}{\oplus}\tilde{\mathfrak{H}}^{1}_{h,0})^{\perp}=(\tilde{\mathfrak{Z}}^{1}_{h,0})^{\perp}$ (87)
+where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (70)–(71).
 4.5 Solving a simpler discrete curl-curl problem
@@
 Since the relations (90) characterize a unique field in $\tilde{V}^{1}_{h,0}$, this will show the existence and uniqueness of a solution.
-The fact that it coincides with the solution of (74) will follow from the fact that the latter satisfies the same relations, as seen in the proof of Theorem 4.2.
+The fact that it coincides with the solution of (84) will follow from the fact that the latter satisfies the same relations, as seen in the proof of Theorem 4.2 and Remark 4.4.
 So, let $\psi_{h}\in\tilde{V}^{0}_{h,\psi}$ be such that ${\bm{p}}^{\psi}_{h,i}=\operatorname{{\bf grad}}\psi_{h}$, and after observing that $\operatorname{{\bf grad}}\psi_{h}\in\tilde{V}^{1}_{h,0}$, take ${\bm{v}}=\operatorname{{\bf grad}}\sigma^{\psi}_{h,i}+{\bm{p}}^{\psi}_{h,i}=\operatorname{{\bf grad}}(\sigma^{\psi}_{h,i}+\psi_{h})$ in (88): since $\bm{v}$} is curl-free, this yields $\operatorname{{\bf grad}}(\sigma^{\psi}_{h,i}+\psi_{h})=0$, hence $\sigma^{\psi}_{h,i}+\psi_{h}=0$ given that this function must vanish on $S_{0}$.